### v1

# Pose2Gaze: Generating Realistic Human Gaze Behaviour from Full-body Poses using an Eye-body Coordination Model

Abstract
While generating realistic body movements, e.g., for avatars in virtual reality, is widely studied in computer vision and graphics, the generation of eye movements that exhibit realistic coordination with the body remains under-explored. We first report a comprehensive analysis of the coordination of human eye and full-body movements during everyday activities based on data from the MoGaze and GIMO datasets. We show that eye gaze has strong correlations with head directions and also full-body motions and there exists a noticeable time delay between body and eye movements. Inspired by the analyses, we then present Pose2Gaze – a novel eye-body coordination model that first uses a convolutional neural network and a spatio-temporal graph convolutional neural network to extract features from head directions and full-body poses respectively and then applies a convolutional neural network to generate realistic eye movements. We compare our method with state-of-the-art methods that predict eye gaze only from head movements for three different generation tasks and demonstrate that Pose2Gaze significantly outperforms these baselines on both datasets with an average improvement of $26.4\%$ and $21.6\%$ in mean angular error, respectively. Our findings underline the significant potential of cross-modal human gaze behaviour analysis and modelling.
[…]
1 Introduction
With the increasing use of augmented and virtual reality (VR)[36], the need to create virtual agents that behave realistically has grown significantly in recent years. Virtual agents that behave like humans is crucial for a variety of applications, such as online learning [31, 32], virtual interviewing and counselling [2, 7], virtual social interactions [40, 29], as well as large-scale virtual worlds [36]. Creating realistic behaviour is also important in character animation, e.g., in films [13]. However, previous work has mainly focused on generating realistic body movements [3, 28] despite the fact that eye movements are equally, if not even more, important in human-human encounters [41] and, as such, also for reducing the uncanny valley effect [42].
Generating realistic eye movement behaviour is challenging given the highly variable spatio-temporal dynamics of gaze allocation as well as various top-down influences, e.g., of the visual world, the agent’s tasks or goals, or social norms that agents have ideally adhere to during interactions with humans [5]. Research on eye gaze prediction has revealed that eye movements are carefully coordinated with head movements and, as such, that head orientation can be used as a proxy to eye gaze direction [45, 21, 20, 18]. Consequently, previous works have mainly focused on exploiting eye-head coordination for generating eye movements [21, 20, 18].
In this work, we provide a complementary perspective and study, for the first time, eye-body coordination during everyday activities and explore the generation of realistic human gaze behaviour from full-body poses.
To this end we first report comprehensive analyses on the correlations between gaze direction and full-body movements on two publicly available datasets: MoGaze [26] and GIMO [51]. These datasets contain human eye and full-body movement data recorded using a body tracking system and a mobile eye tracker during various daily activities.To the best of our knowledge, we are first to study the coordination of eye movements and full-body movements simultaneously. Our analyses confirm that eye gaze direction is strongly correlated with head direction but also with the motions of different human body joints. We also find that there exists a noticeable time delay between full-body motions and eye movements.
Based on these findings we present Pose2Gaze – the first method to generate realistic human gaze behaviour from human full-body poses. At the core of our method is a novel learning-based eye-body coordination model that first uses a convolutional neural network (CNN) and a spatio-temporal graph convolutional neural network (GCN) to extract features from head directions and full-body poses respectively and then employs a convolutional neural network to generate realistic human eye gaze from the extracted features. We compare our method with state-of-the-art gaze prediction methods for three different generation tasks, i.e. generating target eye gaze from past, present, and future body poses respectively, on the MoGaze and GIMO datasets and show that our method outperforms these baselines by a large margin on both the MoGaze ($26.4\%$ improvement) and GIMO datasets ($21.6\%$) in terms of mean angular gaze error [^1] Source code and trained models will be released upon acceptance..
The specific contributions of our work are three-fold:
  • We provide a comprehensive analysis of eye-body coordination in various daily activities and reveal that eye gaze is strongly correlated with head directions and also full-body motions.
  • We propose Pose2Gaze, a novel eye-body coordination model that applies a convolutional neural network and a spatio-temporal graph convolutional neural network to extract features from head directions and full-body poses respectively and uses a convolutional neural network to generate realistic eye gaze.
  • We report extensive experiments on two public datasets for three different generation tasks and demonstrate significant performance improvements over several state-of-the-art methods.
2 Related Work
2.1 Human Movement Generation
Generating realistic human movements for virtual agents is an important research topic in the area of virtual reality. Some researchers focused on generating human body movements from speech signals. For example, Hasegawa et al. used recurrent neural networks (RNNs) to generate natural human full-body poses from perceptual features extracted from the input speech audio[15] while Kucherenko et al. employed autoencoder neural networks to learn latent representations for human poses and speech signals and then learned the mappings between the two representations to generate human body motions[27]. Other researchers used the text transcripts of speech to generate human poses. Specifically, Yoon et al. proposed to use a RNN-based encoder to extract features from speech text and a RNN-based decoder to generate human poses[50] while Bhattacharya et al. employed an end-to-end transformer network to generate human movements from the text transcripts of speech[3]. In addition, there also exist some works that generated human dance motions from input music and achieved good performances[49, 30]. While previous works typically focused on generating human body movements, they neglected the generation of human eye gaze, which is significant for human-human[16, 8] and human-computer interactions[9, 44, 23]. In this work, we focus on generating realistic human eye gaze directly from human full-body poses to fill this gap.
2.2 Eye-body Coordination
The coordination of human eye and body movements is an important research topic in the areas of cognitive science and human-centred computing. Many researchers focused on the coordinated movements between the eyes and the head[21, 46, 12]. Specifically, Stahl analysed the process of gaze shifts and found its amplitude to be proportional to that of head movements[46]. Fang et al. further examined eye-head coordination during visual search in a large visual field and found that eye-head coordination plays an important role in visual cognitive processing[12]. Hu et al. revealed that human eye movements are strongly correlated with head movements in both free-viewing[21, 20] and task-oriented settings[18, 19] in immersive virtual environments while Kothari et al. identified the coordination of eye and head movements in real-world daily activities[24]. Recently, some researchers went beyond eye-head coordination to investigate the correlations between eye movements and the movements of different body parts. For example, Sidenmark et al. focused on the gaze shift process in immersive virtual reality and discovered general eye, head, and torso coordination patterns[43]. Batmaz et al. developed a VR training system based on the coordination of eye and hand movements and compared user performance in three different virtual environments[1]. Emery et al. collected a large-scale dataset that contains human eye, hand, and head movements in a virtual environment and identified the coordination of eye, hand, and head motions[11]. Randhavane et al. investigated the effectiveness of eye-gait coordination on expressing emotions and further employed gaze and gait features to generate various emotions for virtual agents[39]. However, prior works typically focused on the correlations between eye gaze and a particular body part, e.g. head, hand, or torso. In contrast, we are the first to study the coordinations of eye and full-body movements simultaneously, which paves the way for generating realistic eye gaze from full-body poses.
2.3 Eye Gaze Prediction
Eye gaze prediction is a popular research topic in the areas of computer vision and human-centred computing. Typical gaze prediction methods can be classified into bottom-up and top-down approaches. Bottom-up methods predict eye gaze from the low-level image features of the scene content such as intensity, colour, and orientation[22, 6]. For example, Itti et al. used multiscale colour, intensity, and orientation features extracted from the image to predict saliency map (density map of eye gaze distribution)[22] while Cheng et al. employed the global contrast features of the image to generate saliency map[6]. Top-down approaches take high-level features of the scene such as specific tasks and context information into consideration to predict eye gaze[4, 25]. For example, Borji et al. employed players’ input such as 2D mouse positions and joystick buttons to predict their eye gaze[4] while Koulieris et al. used game state variables to predict users’ eye gaze positions in video games[25]. In addition to the typical bottom-up and top-down approaches, recently some researchers took the eye-head coordination into consideration and attempted to predict eye gaze from human head movements[45, 21]. Specifically, Nakashima et al. proposed to use head directions as prior knowledge to improve the accuracy of bottom-up saliency prediction methods through simple multiplication of the predicted saliency map by a Gaussian head direction bias[38]. Sitzmann et al. employed users’ head orientations to predict saliency maps for 360-degree images and achieved an accuracy that is on par with the performance of bottom-up saliency predictors[45]. Hu et al. proposed to use users’ head rotation velocities to predict users’ eye gaze positions in immersive virtual environments and achieved good performances in both free-viewing[21, 20] and task-oriented situations[18, 17]. However, existing works have only explored the effectiveness of head movements on the task of eye gaze prediction. In stark contrast, in this work we focus on the coordination of eye and full-body movements and investigate the effectiveness of full-body movements on generating realistic eye gaze.
3 Analysis of Eye-body Coordination
3.1 Gaze and Motion Data
To study the coordination between human eye gaze and full-body movements, we conducted a comprehensive analysis based on the only public motion capture datasets that also offer eye gaze information: MoGaze [26] and GIMO [51]. The MoGaze dataset contains eye gaze and full-body poses of $21$ human joints from six people performing everyday pick and place activities in an indoor environment. GIMO dataset offers eye gaze and full-body poses of $23$ human joints from $11$ people performing daily activities in various indoor scenes. The activities include resting (sitting or laying on objects), interacting with objects (touching, holding, stepping on, reaching to objects), and changing the state of objects (opening, pushing, transferring, throwing, picking up, lifting, connecting, screwing, grabbing, swapping objects).
3.2 Correlation between Eye Gaze and Body Orientations
We used the forward directions of human body joints available in the datasets to represent body orientations. Specifically, we represented body orientations in MoGaze using the forward directions of head, neck, torso, pelvis, and base and indicated body orientations in GIMO using the directions of head, neck, spine3, spine2, spine1, and pelvis. We used the cosine similarity to analyse the correlation between eye gaze and body orientations. Cosine similarity measures the similarity between two non-zero vectors by calculating the cosine of the angle between the vectors and produces a value in the range from $-1$ to $+1$, where $-1$ indicates perfect negative correlation, 00 denotes no correlation, and $+1$ represents perfect positive correlation. Table 1 summarises the cosine similarities between eye gaze and body orientations in the MoGaze and GIMO datasets. As the table shows, in both datasets eye gaze direction is strongly correlated with the directions of different body joints. In particular, eye gaze exhibits very high correlation with head direction, achieving a cosine similarity of $0.92$ in MoGaze and $0.93$ in GIMO. We also find that the directions of different body joints have very strong correlations with each other, e.g. neck and head directions have a cosine similarity of $0.95$ in MoGaze and $0.92$ in GIMO, indicating that there exists redundant information among these directions. Therefore, for simplicity, we suggest to use only head direction to generate human eye gaze rather than using all the body directions.
Gaze Head Neck Torso Pelvis Base
Gaze 1 0.92 0.84 0.64 0.59 0.63
Head - 1 0.95 0.81 0.78 0.81
Neck - - 1 0.94 0.91 0.93
Torso - - - 1 0.99 0.99
Pelvis - - - - 1 0.99
Base - - - - - 1
(a) Gaze-body orientations in MoGaze
Gaze Head Neck Spine3 Spine2 Spine1 Pelvis
Gaze 1 0.93 0.89 0.84 0.77 0.84 0.72
Head - 1 0.92 0.84 0.77 0.84 0.71
Neck - - 1 0.97 0.95 0.97 0.90
Spine3 - - - 1 0.98 0.99 0.94
Spine2 - - - - 1 0.97 0.97
Spine1 - - - - - 1 0.93
Pelvis - - - - - - 1
(b) Gaze-body orientations in GIMO
Table 1: The cosine similarities between eye gaze directions and the directions of different body joints in the MoGaze and GIMO datasets. Gaze directions are strongly correlated with body orientations, especially with head directions. The directions of different body joints have strong correlations with each other.
To investigate whether there exist time delays between head direction and eye gaze, we added different time intervals between head and gaze directions and further calculated their cosine similarities. We can see from Figure 1 that head direction achieves its highest correlation with eye gaze at $100$ $ms$ in MoGaze and 00 $ms$ in GIMO, indicating that there exists little or no time delay between head and eye movements.
(a) Gaze-head direction in MoGaze
(b) Gaze-head direction in GIMO
Figure 1: The cosine similarities between head and gaze directions at different time intervals in the (a) MoGaze and (b) GIMO datasets. Time interval refers to the time difference between head and gaze directions. The highest gaze-head correlation occurs at $100$ $ms$ in MoGaze and 00 $ms$ in GIMO, suggesting that there is little or no time delay between head and eye movements.
3.3 Correlations between Eye Gaze and Body Movements
To analyse the correlations between human eye gaze and human full-body movements, we calculated the velocities of different body joints and normalised these velocities to 3D unit vectors to represent the directions of body motions. We further calculated the cosine similarities between eye gaze and the directions of body motions to measure their correlations. Figure 2 shows the cosine similarities between eye gaze and full-body motions in the MoGaze and GIMO datasets, respectively. We can see that eye gaze directions have strong correlations with the motions of different body joints in both datasets.
(a) Gaze-body motions in MoGaze
(b) Gaze-body motions in GIMO
Figure 2: The cosine similarities between eye gaze directions and the directions of body motions in the (a) MoGaze and (b) GIMO datasets. Eye gaze is strongly correlated with human full-body motions.
To investigate any potential time delays between body motions and eye movements, we added different time intervals between body motions and gaze directions and further calculated their correlations. Specifically, for simplicity we first grouped all the human joints into three different body parts, i.e. Torso, Arm, and Leg. In the MoGaze dataset Torso consists of the body joints of base, pelvis, torso, neck, and head while in the GIMO dataset Torso contains the body joints of pelvis, spine1, spine2, spine3, neck, head, and jaw. In both datasets, Arm contains the body joints of collar, shoulder, elbow, and wrist while Leg includes hip, knee, ankle, and foot, where the left and right joints are averaged into one single joint, e.g. left shoulder and right shoulder are represented using shoulder. We then calculated the cosine similarities between eye gaze and the motions of different body parts as well as between eye gaze and the motions of specific joints. The correlation between gaze and a body part was computed using the mean of the cosine similarities between eye gaze and the specific joints within the body part. We can see from Figure 3 that in both grouped body parts and specific body joints, the highest correlations consistently occur between current eye gaze and body motions in the near future, suggesting that there exists a noticeable time delay between body motions and eye movements.
(a) Gaze-body motions in MoGaze
(b) Gaze-torso motions in MoGaze
(c) Gaze-arm motions in MoGaze
(d) Gaze-leg motions in MoGaze
(e) Gaze-body motions in GIMO
(f) Gaze-torso motions in GIMO
(g) Gaze-arm motions in GIMO
(h) Gaze-leg motions in GIMO
Figure 3: The cosine similarities between eye gaze and the motions of different body parts as well as between eye gaze and the motions of specific body joints at different time intervals in the (a-d) MoGaze and (e-h) GIMO datasets. Time interval refers to the time difference between body movements and eye gaze. The highest correlations consistently occur between current eye gaze and body motions in the near future (Time Interval $>0$), indicating that there exists a significant time delay between body and eye movements.
3.4 Summary
Through comprehensive analysis, we find that eye gaze has very strong correlations with body orientations, especially with head directions, and there exists little or no time delay between head and eye movements. We also find that eye gaze is strongly correlated with human full-body motions and there exists a noticeable time delay between body and eye movements. These results suggest that head directions and full-body poses contain rich information about eye gaze and thus can be used to generate realistic eye movements.
4 Pose2Gaze: Eye-body Coordination Model
We define pose-based eye gaze generation as the task of generating a sequence of eye gaze directions $G_{t+1:t+T}=\{g_{t+1},g_{t+2},...,g_{t+T}\}\in R^{3\times T}$, where $g$ is a 3D unit vector and $T$ is the length of the target eye gaze sequence, from human body orientations and motions. We use a sequence of head directions $H_{t+1+\Delta t_{h}:t+T+\Delta t_{h}}=\{h_{t+1+\Delta t_{h}},h_{t+2+\Delta t_{h}},...,h_{t+T+\Delta t_{h}}\}\in R^{3\times T}$ to represent body orientations, where $h$ is a 3D unit vector and $\Delta t_{h}$ is the time interval between the input head directions and target eye gaze. We employ a sequence of the 3D positions of all human joints $P_{t+1+\Delta t_{p}:t+T+\Delta t_{p}}=\{p_{t+1+\Delta t_{p}},p_{t+2+\Delta t_{p}},...,p_{t+T+\Delta t_{p}}\}\in R^{3\times N\times T}$ to represent body motions, where $N$ is the number of human joints and $\Delta t_{p}$ is the time interval between the input body motions and target eye gaze. By setting different time intervals between the input body movements and the target gaze directions, our model can be trained to generate target eye gaze from past, present, or future body poses, respectively (see subsection 5.1). Our method consists of three main components: a body orientation feature extraction module that extracts orientation features from head directions, a body motion feature extraction module that extracts motion features from full-body poses, and an eye gaze generation module that generates gaze directions from the extracted body orientation and motion features (see Figure 4 for an overview of our method).
Figure 4: Architecture of the proposed Pose2Gaze model. Pose2Gaze first uses a 1D convolutional neural network to extract body orientation features from head directions, then applies a spatio-temporal graph convolutional neural network to extract the body motion features from human full-body poses, and finally employs a 1D convolutional neural network to generate human eye gaze from the extracted body orientation and motion features.
4.1 Body Orientation Feature Extraction
In light of the good performance of 1D convolutional neural networks for processing time series data [20, 18], we employed three 1D CNN layers to extract body orientation features from the sequence of head directions. Specifically, we used two 1D CNN layers, each with $32$ channels and a kernel size of three, to process the head direction sequence $H\in R^{3\times T}$. Each CNN layer was followed by a layer normalisation (LN) and a Tanh activation function. After the two CNN layers, we used a 1D CNN layer with $32$ channels and a kernel size of three, and a Tanh activation function to obtain the body orientation features $f_{ori}\in R^{32\times T}$.
4.2 Body Motion Feature Extraction
Given the effectiveness of discrete cosine transform (DCT) for extracting temporal features from human pose data [33, 35], we first employed DCT to encode human pose $P\in R^{3\times N\times T}$ in the temporal domain using DCT matrix $M_{dct}\in R^{T\times T}$:
$P_{dct}=PM_{dct},$ (1)
where $P_{dct}\in R^{3\times N\times T}$ is the human pose after DCT transform. Considering the good performance of graph convolutional neural networks for processing human pose data [33, 35], we propose two GCN blocks, i.e. a start GCN block and a residual GCN block, to extract motion features from the transformed pose data.
Start GCN Block
The start GCN block first applies a temporal GCN (T-GCN) to extract the temporal features from the transformed pose data $P_{dct}$. The temporal GCN views the pose data as a fully-connected graph that contains $T$ nodes corresponding to pose data at $T$ time steps. It learns the weighted adjacency matrix $A^{T}\in R^{T\times T}$ of the fully-connected temporal graph and performs temporal convolution using
$f_{temp}=P_{dct}A^{T},$ (2)
where $f_{temp}\in R^{3\times N\times T}$ is the extracted temporal features. $f_{temp}$ was then permuted to $f_{temp}\in R^{T\times N\times 3}$. A weight matrix $W^{start}\in R^{3\times 16}$ was applied to convert the input node features ($3$ dimensions) to latent features ($16$ dimensions):
$f_{lat}=f_{temp}W^{start},$ (3)
where $f_{lat}\in R^{T\times N\times 16}$ is the latent features. After the weight matrix, a spatial GCN (S-GCN) was applied to extract the spatial features. The spatial GCN views the latent features $f_{lat}$ as a fully-connected graph that contains $N$ nodes corresponding to $N$ human joints. S-GCN learns the weighted adjacency matrix $A^{S}\in R^{N\times N}$ of the fully-connected spatial graph and performs spatial convolution using
$f_{spa}=A^{S}f_{lat},$ (4)
where $f_{spa}\in R^{T\times N\times 16}$ is the extracted spatial features. $f_{spa}$ was further permuted to $f_{spa}\in R^{16\times N\times T}$. The output of the spatial GCN is copied along the temporal dimension ($R^{16\times N\times T}\rightarrow R^{16\times N\times 2T}$) to enhance the features [33] and is then used as input to the residual GCN block.
Residual GCN Block
The residual GCN block contains $m$ GCN components with each component consisting of a temporal GCN that learns the temporal adjacency matrix $A^{T}_{i}\in R^{2T\times 2T}$, a weight matrix $W^{res}_{i}\in R^{16\times 16}$ that extracts the latent features, a spatial GCN that learns the spatial adjacency matrix $A^{S}_{i}\in R^{N\times N}$, a layer normalisation, a Tanh activation function, and a dropout layer with dropout rate $0.3$ to avoid overfitting. We set the number of GCN components $m$ to $4$ and added a residual connection for each GCN component to improve the network flow. We further cut the output of the residual GCN block in half in the temporal dimension to reduce the feature dimensions and obtained the spatio-temporal human body motion features $f_{mot}\in R^{16\times N\times T}$.
4.3 Eye Gaze Generation
To generate eye gaze from the extracted body orientation and motion features, we first aggregated the body motion features $f_{mot}$ along the spatial dimension, i.e. concatenated the features of different body joints into a single motion feature ($R^{16\times N\times T}\rightarrow R^{16N\times T}$). We then fused the extracted orientation and motion features by concatenating them along the spatial dimension and obtained $f\in R^{(16N+32)\times T}$. We finally used a 1D convolutional neural network to generate eye gaze from the fused features. Specifically, we used two CNN layers, each with a kernel size of three, to process the fused features. The first CNN layer has $64$ channels and is followed by a layer normalisation and a Tanh activation function while the second CNN layer uses three channels and a Tanh activation function to generate the eye gaze. The generated eye gaze directions were finally normalised to unit vectors: $\hat{G}_{t+1:t+T}=\{\hat{g}_{t+1},\hat{g}_{t+2},...,\hat{g}_{t+T}\}\in R^{3\times T}$.
4.4 Loss Function
To train our model, we used the mean angular error between the generated eye gaze directions $\hat{G}_{t+1:t+T}$ and the ground truth gaze directions $G_{t+1:t+T}$ as our loss function:
$\ell=\frac{1}{T}\sum_{j=t+1}^{t+T}\arccos(\hat{g_{j}}\cdot g_{j}).$ (5)
5 Experiments and Results
We conducted extensive experiments to evaluate the performance of our gaze generation model. Specifically, we compared our model with state-of-the-art gaze prediction methods that estimate eye gaze from head movements on both the MoGaze and GIMO datasets. We tested these methods under three different generation settings, i.e. generating target eye gaze from past, present, and future body poses, respectively. We also performed an ablation study to evaluate the effectiveness of each component used in our model.
5.1 Experimental Settings
Datasets
We evaluated our method on both the MoGaze and GIMO datasets since they are the two most established public datasets that offer both eye gaze and human full-body poses. MoGaze contains eye gaze and full-body poses recorded at $30$ fps from six people performing everyday pick and place actions. To evaluate our model’s generalisation capability for different users, we performed a leave-one-person-out cross-validation: We trained on the data from five people and tested on the remaining one, repeated this procedure six times by testing for a different target person, and calculated the average performance across all six iterations. GIMO contains eye gaze and pose data ($30$ fps) from $11$ people performing daily activities in various indoor scenes. To evaluate our model on GIMO, we used the default training and test sets provided by the authors [51], i.e. we used data from $12$ scenes for training and data from $14$ scenes ($12$ known scenes and two new scenes) for testing.
Evaluation Metric
As is commonly used in gaze prediction [20, 18], we employed the mean angular error between the generated and ground truth gaze directions (see Equation 5) as the metric to evaluate model performance.
Baselines
We compared our method with the following state-of-the-art gaze prediction methods that generate eye gaze from human head movements:
  • Head Direction: Head Direction has been frequently used as a proxy for eye gaze due to the strong link between eye and head movements [45, 20, 38].
  • DGaze [20]: DGaze predicts eye gaze from the sequence of head movements using a 1D convolutional neural network.
  • FixationNet [18]: FixationNet extracts features from head movement sequence using a 1D convolutional neural network and combines the features with prior knowledge of gaze distribution to generate eye gaze directions.
Implementation Details
We trained the baseline methods from scratch using their default parameters. To train our method, we used the Adam optimiser with an initial learning rate of $0.005$ that we decayed by $0.95$ every epoch. We used a batch size of $32$ to train our method for a total of $50$ epochs. We implemented our method using the PyTorch framework.
Generation Settings
We set our model to generate $15$ frames (corresponding to $500$ ms) of eye gaze directions $G_{t+1:t+15}=\{g_{t+1},g_{t+2},...,g_{t+15}\}$. We evaluated our method for three different generation tasks: Generating target eye gaze from past, present, and future body poses, respectively:
  • Generating Gaze from Past Poses: We used the body poses and head directions in the past $15$ frames $P_{t-14:t}=\{p_{t-14},p_{t-13},...,p_{t}\}$ and $H_{t-14:t}=\{h_{t-14},h_{t-13},...,h_{t}\}$ as input to generate the target eye gaze. This setting is equivalent to predicting human eye gaze in the future (gaze forecasting) [18, 20], which is important for a variety of applications including visual attention enhancement [10], dynamic event triggering [14], as well as human-human and human-computer interaction [47, 37].
  • Generating Gaze from Present Poses: We used the body poses and head directions at the present time $P_{t+1:t+15}=\{p_{t+1},p_{t+2},...,p_{t+15}\}$ and $H_{t+1:t+15}=\{h_{t+1},h_{t+2},...,h_{t+15}\}$ as input to generate the corresponding gaze directions. This setting corresponds to estimating human eye gaze in real time [21, 25], which is key for a number of applications including gaze-contingent rendering [21, 20], gaze-based interaction [34], and gaze-guided redirected walking [48].
  • Generating Gaze from Future Poses: subsection 3.3 reveals that eye gaze has highest correlations with body motions in the near future and this implies that using future body poses may improve the performance of gaze generation. To evaluate the effectiveness of future body poses, we used the body poses in the future $15$ frames $P_{t+16:t+30}=\{p_{t+16},p_{t+17},...,p_{t+30}\}$ and the head directions at the present time $H_{t+1:t+15}=\{h_{t+1},h_{t+2},...,h_{t+15}\}$ as input to generate the target eye gaze. We used real-time head directions because there exists little or no time delay between head and eye movements (subsection 3.2). This setting can be seen as an offline processing way of generating realistic eye movements and is particularly important for creating and animating virtual humans [3, 41].
5.2 Gaze Generation Results
Generating Gaze from Past Poses
Table 2 summarises the performances of different methods for generating eye gaze from past poses on both datasets. When evaluating on the MoGaze dataset, we calculated the mean angular errors on individual actions (MoGaze-pick, MoGaze-place) as well as on all the actions (MoGaze-all). We can see from the table that our method consistently outperforms the state-of-the-art methods on the MoGaze dataset. For pick and place actions, our method achieves an average improvement of $18.7\%$ ($14.8^{\circ}$ vs. $18.2^{\circ}$) and $27.0\%$ ($11.1^{\circ}$ vs. $15.2^{\circ}$) respectively in terms of mean angular error. For all actions on MoGaze, our method achieves an average improvement of $22.0\%$ ($13.1^{\circ}$ vs. $16.8^{\circ}$) over the state of the art. The action labels for some recordings in the GIMO dataset are missing. Therefore, when evaluating on GIMO, we directly calculated the mean angular errors on all the actions (GIMO-all) and indicated the results in Table 2. We find that our method significantly outperforms the state of the art, achieving an average improvement of $9.7\%$ ($18.6^{\circ}$ vs. $20.6^{\circ}$). We further performed a paired Wilcoxon signed-rank test to compare the performances of our method with the state-of-the-art methods and validated that the differences between our method and the state of the art are statistically significant ($p<0.01$). The above results demonstrate that our method has a strong capability of generating eye gaze from past body poses.
Ours FixationNet[18] DGaze[20] Head Direction
MoGaze-pick $\textbf{14.8}^{\circ}$ $\textit{18.2}^{\circ}$ $18.3^{\circ}$ $37.8^{\circ}$
MoGaze-place $\textbf{11.1}^{\circ}$ $\textit{15.2}^{\circ}$ $15.3^{\circ}$ $34.9^{\circ}$
MoGaze-all $\textbf{13.1}^{\circ}$ $\textit{16.8}^{\circ}$ $16.9^{\circ}$ $36.4^{\circ}$
GIMO-all $\textbf{18.6}^{\circ}$ $21.1^{\circ}$ $\textit{20.6}^{\circ}$ $23.1^{\circ}$
Table 2: Mean angular errors of different methods for generating eye gaze from past poses on the MoGaze and GIMO datasets. Best results are in bold while the second best are in italic.
Generating Gaze from Present Poses
Table 3 summarises the mean angular errors of different methods for generating eye gaze from present body poses. We can see that on the MoGaze dataset our method significantly outperforms the state of the art on individual actions (MoGaze-pick, MoGaze-place) as well as on all the actions (MoGaze-all). Specifically, for pick, place and all actions, our method achieves an average improvement of $20.5\%$ ($10.5^{\circ}$ vs. $13.2^{\circ}$), $20.5\%$ ($9.3^{\circ}$ vs. $11.7^{\circ}$), and $20.8\%$ ($9.9^{\circ}$ vs. $12.5^{\circ}$) respectively in terms of mean angular error. On the GIMO dataset, our method outperforms the state of the art with an average improvement of $17.5\%$ ($16.0^{\circ}$ vs. $19.4^{\circ}$). A paired Wilcoxon signed-rank test was performed to compare the performances of our method with the state-of-the-art methods and the results indicated that the differences between our method and the state of the art are statistically significant ($p<0.01$). Figure 5 illustrates the generation results of different methods on the MoGaze and GIMO datasets. We can see that our method exhibits better generation performance than other methods (See supplementary video for more results). These results demonstrate that our method is able to generate realistic eye gaze from present body poses.
Ours FixationNet[18] DGaze[20] Head Direction
MoGaze-pick $\textbf{10.5}^{\circ}$ $\textit{13.2}^{\circ}$ $13.4^{\circ}$ $17.6^{\circ}$
MoGaze-place $\textbf{9.3}^{\circ}$ $\textit{11.7}^{\circ}$ $12.1^{\circ}$ $16.2^{\circ}$
MoGaze-all $\textbf{9.9}^{\circ}$ $\textit{12.5}^{\circ}$ $12.8^{\circ}$ $16.9^{\circ}$
GIMO-all $\textbf{16.0}^{\circ}$ $19.5^{\circ}$ $\textit{19.4}^{\circ}$ $19.4^{\circ}$
Table 3: Mean angular errors of different methods for generating eye gaze from present poses on the MoGaze and GIMO datasets.
Figure 5: Results of different methods for generating eye gaze from present poses on MoGaze and GIMO. The green line indicates the ground truth while the blue line represents the generated eye gaze.
Generating Gaze from Future Poses
Table 4 summarises the mean angular errors of different methods for generating eye gaze from future body poses. We can see that for pick, place and all actions on the MoGaze dataset, our method consistently outperforms the state of the art, achieving an average improvement of $25.8\%$ ($9.8^{\circ}$ vs. $13.2^{\circ}$), $26.5\%$ ($8.6^{\circ}$ vs. $11.7^{\circ}$), and $26.4\%$ ($9.2^{\circ}$ vs. $12.5^{\circ}$) respectively. On the GIMO dataset, our method outperforms the state of the art with an average improvement of $21.6\%$ ($15.2^{\circ}$ vs. $19.4^{\circ}$) in terms of mean angular error. A paired Wilcoxon signed-rank test was conducted and the results validated that the differences between our method and the state of the art are statistically significant ($p<0.01$). The above results demonstrate that our method has high performance in generating eye gaze from future body poses.
Ours FixationNet[18] DGaze[20] Head Direction
MoGaze-pick $\textbf{9.8}^{\circ}$ $\textit{13.2}^{\circ}$ $13.4^{\circ}$ $17.6^{\circ}$
MoGaze-place $\textbf{8.6}^{\circ}$ $\textit{11.7}^{\circ}$ $12.1^{\circ}$ $16.2^{\circ}$
MoGaze-all $\textbf{9.2}^{\circ}$ $\textit{12.5}^{\circ}$ $12.8^{\circ}$ $16.9^{\circ}$
GIMO-all $\textbf{15.2}^{\circ}$ $19.5^{\circ}$ $\textit{19.4}^{\circ}$ $19.4^{\circ}$
Table 4: Mean angular errors of different methods for generating eye gaze from future poses on the MoGaze and GIMO datasets.
5.3 Summary
The results in above section demonstrate that our method significantly outperforms the state-of-the-art methods for three different generation tasks. Furthermore, by comparing our method’s performances under different generation settings, we find that our method achieves its best performance in using future body poses ($9.2^{\circ}$ on MoGaze and $15.2^{\circ}$ on GIMO), followed by using present body poses ($9.9^{\circ}$ on MoGaze and $16.0^{\circ}$ on GIMO) and then using past poses ($13.1^{\circ}$ on MoGaze and $18.6^{\circ}$ on GIMO). This result corresponds with our analysis in subsection 3.3 that eye gaze has highest correlations with body motions in the near future and suggests that in offline applications where future pose is available, e.g. creating and animating virtual humans [3, 41], using future body poses could generate more realistic eye gaze.
Our model has smaller size than the state-of-the-art methods, containing only $0.17M$ trainable parameters while DGaze has $0.27M$ parameters and FixationNet contains $0.37M$ parameters. We implemented our model on an NVIDIA Tesla V100 SXM2 32GB GPU with an Intel(R) Xeon(R) Platinum 8260 CPU @ 2.40GHz and calculated its time costs. We find that our model requires only $35$ ms to train per batch and $3$ ms to test per batch, suggesting that our model is fast enough for practical usage.
5.4 Ablation Study
We finally performed an ablation study to evaluate the effectiveness of each component used in our model.
Effectiveness of Our Model Architecture
To evaluate the effectiveness of our model architecture, we evaluated different ablated versions of it that did not contain the DCT, spatial GCN, temporal GCN, the body poses input, and the head directions input, and re-trained the ablated models. Table 5 summarises the performances of different ablated versions of our method for generating eye gaze on the GIMO dataset. We can see that our method consistently outperforms the ablated methods under different generation settings and the results are statistically significant (paired Wilcoxon signed-rank test, $p<0.01$), thus underlining the effectiveness of our model architecture. One exception is that our method is not as good as the w/o DCT and w/o T-GCN versions in generating eye gaze from past body poses. This is probably because past body poses have weaker correlations with eye gaze than present and future body poses (subsection 3.3), and thus the temporal features extracted from past body poses using DCT and temporal GCN are not as effective as that extracted from present and future poses.
Ours w/o DCT w/o S-GCN w/o T-GCN w/o Poses w/o Head
Past $18.6^{\circ}$ $\textbf{18.2}^{\circ}$ $19.0^{\circ}$ $18.3^{\circ}$ $22.8^{\circ}$ $23.2^{\circ}$
Present $\textbf{16.0}^{\circ}$ $16.6^{\circ}$ $17.3^{\circ}$ $16.1^{\circ}$ $20.7^{\circ}$ $20.4^{\circ}$
Future $\textbf{15.2}^{\circ}$ $16.4^{\circ}$ $17.2^{\circ}$ $16.0^{\circ}$ $20.7^{\circ}$ $19.5^{\circ}$
Table 5: Mean angular errors of different ablated versions of our method for generating eye gaze under different settings on GIMO.
Importance of Different Body Parts
We further evaluated the importance of different body parts for eye gaze generation. To this end, we re-trained our model using different body parts as input, i.e. only using the Torso part, Arm part, Leg part, as well as different combinations of these. Table 6 shows the performances of our method for generating eye gaze from different body parts on the GIMO dataset. As can be seen from the table, our method achieves significantly better performances than other versions (paired Wilcoxon signed-rank test, $p<0.01$), validating that each body part used in our model can help improve the performance of eye gaze generation.
Ours Torso Arm Leg Torso+Arm Torso+Leg Arm+Leg
Past $\textbf{18.6}^{\circ}$ $20.7^{\circ}$ $19.4^{\circ}$ $19.7^{\circ}$ $19.6^{\circ}$ $19.6^{\circ}$ $18.8^{\circ}$
Present $\textbf{16.0}^{\circ}$ $17.8^{\circ}$ $17.4^{\circ}$ $18.0^{\circ}$ $16.2^{\circ}$ $17.0^{\circ}$ $16.9^{\circ}$
Future $\textbf{15.2}^{\circ}$ $18.6^{\circ}$ $16.4^{\circ}$ $18.3^{\circ}$ $16.0^{\circ}$ $16.7^{\circ}$ $16.7^{\circ}$
Table 6: Mean angular errors of our method for generating eye gaze from different body parts under different settings on GIMO.
6 Discussion
In this work we have made an important step towards understanding the correlation between eye and full-body movements in daily activities and generating eye gaze from full-body poses.
Eye-body Coordination
Our analyses on eye-body coordination in daily activities effectively guided the design of our model. Specifically, in the analysis of correlation between eye gaze and body orientations, we find that eye gaze has strong correlations with the orientations of different body joints, especially with head directions, and these body orientations are strongly correlated with each other (Table 1). Inspired by this, our model extracts body orientation features directly from head directions rather than all the body orientations (subsection 4.1) because head direction itself contains sufficient information for generating eye gaze. We also find that there exists little or no time delay between head and eye movements (Figure 1). Therefore, in the task of generating gaze from future poses (subsection 5.1), we used head directions at the present time to generate target eye gaze rather than using future head directions. In the analysis of correlation between eye gaze and body motions, we find that eye gaze is strongly correlated with full-body motions (Figure 2). Therefore, our model extracts features from human full-body poses to generate realistic eye gaze (subsection 4.2). We also find that there exists a noticeable time delay between full-body motions and eye gaze (Figure 3). Inspired by this, we used future body poses as input to improve the performance of eye gaze generation in situations where future pose information is available (subsection 5.2). The good performance of our model validated that our analyses are effective and significant for the design of eye-body coordination models.
Pose-based Gaze Generation
Prior methods have typically generated eye gaze from head movements while our method generates eye gaze from full-body poses. Our method significantly outperforms prior methods on both the MoGaze and GIMO datasets for three different generation tasks (Table 2, Table 3, Table 4) and the ablation studies validated that each body part contributes to the performance of our method (Table 6). These results reveal the significant potential of human full-body poses for generating realistic eye movements and thus open the promising research direction of pose-based gaze generation.
Applications of Our Method
We evaluated our method for three different generation tasks that are respectively linked to different kinds of applications (subsection 5.1). The fact that our method outperforms the state-of-the-art methods by a large margin under all the three different generation settings (Table 2, Table 3, Table 4) demonstrates that our method has great potential for these numerous applications including dynamic event triggering [14], gaze-contingent rendering [21, 20], gaze-based interaction [9], as well as creating and animating virtual humans [3, 41]. In addition, with the development of motion capture and motion generation techniques [3, 15], it becomes increasingly easy to obtain human full-body poses. Our approach could be integrated with such motion capture and motion generation techniques to generate realistic eye and full-body movements simultaneously.
Limitations
Despite these advances, we identified several limitations of our work that we plan to address in the future. First, we conducted our analysis and experiments using two public indoor datasets that contain eye and full-body movements in various daily activities. However, there exist some other interesting activities that are not covered by these datasets, e.g., narration or conversation actions [3]. No suitable gaze dataset exists yet, so the coordination of eye and full-body movements for these activities still remains to be explored. In addition, all activities in our datasets were performed in indoor environments and it remains to be seen whether eye-body coordination as characterised here also applies to outdoor scenarios. Furthermore, the activities we studied only cover the interaction between humans and objects and don’t involve other interaction paradigms, e.g. interaction between two humans and interaction between a human and a social robot. Eye-body coordination in such interaction paradigms may be different and deserves to be studied in future work.
Future Work
Besides overcoming the above limitations, many potential avenues of future work exist. First, it is well-known that human eye gaze is influenced by both the bottom-up scene content and the top-down tasks. Therefore, it will be interesting to analyse how the bottom-up and top-down factors influence the eye-body coordination. In addition, incorporating other modalities such as audio, text, or speech into our model may further boost our model’s gaze generation performance. Finally, generating stylistic eye gaze, e.g. eye gaze that can convey different emotions [39], is an interesting avenue to extend our work.
7 Conclusion
In this work we were the first to explore the challenging task of generating realistic human eye gaze from human full-body poses. We first conducted a comprehensive analysis on the coordination of human eye and full-body movements in everyday activities and revealed that eye gaze is strongly correlated with head directions and also full-body motions. Based on these insights, we proposed a novel eye-body coordination model to generate eye gaze from head directions and full-body poses that outperformed the state-of-the-art methods by a large margin for three different generation tasks. Taken together, our work provides novel insights into eye-body coordination during daily activities and makes an important step towards more holistic, cross-modal generation of virtual agent behaviour.

### v3

@@
-Pose2Gaze: Generating Realistic Human Gaze Behaviour from Full-body Poses using an Eye-body Coordination Model
+Pose2Gaze: Eye-body Coordination during Daily Activities for Gaze Prediction from Full-body Poses
 Abstract
-While generating realistic body movements, e.g., for avatars in virtual reality, is widely studied in computer vision and graphics, the generation of eye movements that exhibit realistic coordination with the body remains under-explored.
-We first report a comprehensive analysis of the coordination of human eye and full-body movements during everyday activities based on data from the MoGaze and GIMO datasets.
-We show that eye gaze has strong correlations with head directions and also full-body motions and there exists a noticeable time delay between body and eye movements.
-Inspired by the analyses, we then present Pose2Gaze – a novel eye-body coordination model that first uses a convolutional neural network and a spatio-temporal graph convolutional neural network to extract features from head directions and full-body poses respectively and then applies a convolutional neural network to generate realistic eye movements.
-We compare our method with state-of-the-art methods that predict eye gaze only from head movements for three different generation tasks and demonstrate that Pose2Gaze significantly outperforms these baselines on both datasets with an average improvement of $26.4\%$ and $21.6\%$ in mean angular error, respectively.
-Our findings underline the significant potential of cross-modal human gaze behaviour analysis and modelling.
-[…]
-[…]
-1 Introduction
-With the increasing use of augmented and virtual reality (VR)[36], the need to create virtual agents that behave realistically has grown significantly in recent years.
-Virtual agents that behave like humans is crucial for a variety of applications, such as online learning [31, 32], virtual interviewing and counselling [2, 7], virtual social interactions [40, 29], as well as large-scale virtual worlds [36].
-Creating realistic behaviour is also important in character animation, e.g., in films [13].
-However, previous work has mainly focused on generating realistic body movements [3, 28] despite the fact that eye movements are equally, if not even more, important in human-human encounters [41] and, as such, also for reducing the uncanny valley effect [42].
-Generating realistic eye movement behaviour is challenging given the highly variable spatio-temporal dynamics of gaze allocation as well as various top-down influences, e.g., of the visual world, the agent’s tasks or goals, or social norms that agents have ideally adhere to during interactions with humans [5].
-Research on eye gaze prediction has revealed that eye movements are carefully coordinated with head movements and, as such, that head orientation can be used as a proxy to eye gaze direction [45, 21, 20, 18].
-Consequently, previous works have mainly focused on exploiting eye-head coordination for generating eye movements [21, 20, 18].
-In this work, we provide a complementary perspective and study, for the first time, eye-body coordination during everyday activities and explore the generation of realistic human gaze behaviour from full-body poses.
-To this end we first report comprehensive analyses on the correlations between gaze direction and full-body movements on two publicly available datasets: MoGaze [26] and GIMO [51].
-These datasets contain human eye and full-body movement data recorded using a body tracking system and a mobile eye tracker during various daily activities.To the best of our knowledge, we are first to study the coordination of eye movements and full-body movements simultaneously.
-Our analyses confirm that eye gaze direction is strongly correlated with head direction but also with the motions of different human body joints.
-We also find that there exists a noticeable time delay between full-body motions and eye movements.
-Based on these findings we present Pose2Gaze – the first method to generate realistic human gaze behaviour from human full-body poses.
-At the core of our method is a novel learning-based eye-body coordination model that first uses a convolutional neural network (CNN) and a spatio-temporal graph convolutional neural network (GCN) to extract features from head directions and full-body poses respectively and then employs a convolutional neural network to generate realistic human eye gaze from the extracted features.
-We compare our method with state-of-the-art gaze prediction methods for three different generation tasks, i.e.
-generating target eye gaze from past, present, and future body poses respectively, on the MoGaze and GIMO datasets and show that our method outperforms these baselines by a large margin on both the MoGaze ($26.4\%$ improvement) and GIMO datasets ($21.6\%$) in terms of mean angular gaze error [^1] Source code and trained models will be released upon acceptance..
-The specific contributions of our work are three-fold:
-  • We provide a comprehensive analysis of eye-body coordination in various daily activities and reveal that eye gaze is strongly correlated with head directions and also full-body motions.
-  • We propose Pose2Gaze, a novel eye-body coordination model that applies a convolutional neural network and a spatio-temporal graph convolutional neural network to extract features from head directions and full-body poses respectively and uses a convolutional neural network to generate realistic eye gaze.
-  • We report extensive experiments on two public datasets for three different generation tasks and demonstrate significant performance improvements over several state-of-the-art methods.
-2 Related Work
-2.1 Human Movement Generation
-Generating realistic human movements for virtual agents is an important research topic in the area of virtual reality.
-Some researchers focused on generating human body movements from speech signals.
+Human eye gaze plays a significant role in many virtual and augmented reality (VR/AR) applications, such as gaze-contingent rendering, gaze-based interaction, or eye-based activity recognition.
+However, prior works on gaze analysis and prediction have only explored eye-head coordination and were limited to human-object interactions.
+We first report a comprehensive analysis of eye-body coordination in various human-object and human-human interaction activities based on four public datasets collected in real-world (MoGaze), VR (ADT), as well as AR (GIMO and EgoBody) environments.
+We show that in human-object interactions, e.g.
+pick and place, eye gaze exhibits strong correlations with full-body motion while in human-human interactions, e.g.
+chat and teach, a person’s gaze direction is correlated with the body orientation towards the interaction partner.
+Informed by these analyses we then present Pose2Gaze – a novel eye-body coordination model that uses a convolutional neural network and a spatio-temporal graph convolutional neural network to extract features from head direction and full-body poses, respectively, and then uses a convolutional neural network to predict eye gaze.
+We compare our method with state-of-the-art methods that predict eye gaze only from head movements and show that Pose2Gaze outperforms these baselines with an average improvement of $24.0\%$ on MoGaze, $10.1\%$ on ADT, $21.3\%$ on GIMO, and $28.6\%$ on EgoBody in mean angular error, respectively.
+We also show that our method significantly outperforms prior methods in the sample downstream task of eye-based activity recognition.
+These results underline the significant information content available in eye-body coordination during daily activities and open up a new direction for gaze prediction.
+Index Terms:
+Eye-body coordination, human-object interaction, human-human interaction, gaze prediction, activity recognition, virtual reality, augmented reality
+I Introduction
+With the increasing use of virtual and augmented reality (VR/AR), understanding and predicting human behaviour in VR/AR environments has become a popular research topic and is a key requirement for intelligent human-aware VR/AR agents [1, 2].
+Human eye movements are of particular interest given that they are key for important VR/AR applications, such as 1) gaze-contingent rendering that renders the gaze central region with high quality while decreasing the fidelity in peripheral region to increase rendering efficiency [3]; 2) gaze-based interaction that uses eye gaze to interact with 3D objects in virtual environments [4]; 3) virtual content optimisation that optimises the design of virtual content based on gaze distributions [5]; 4) gaze-guided redirected walking that redirects a user’s walking path based on their gaze directions [6]; or 5) eye-based activity recognition that recognises user activities from their eye movements [7].
+Understanding and predicting human eye movements in VR/AR is challenging given the highly variable spatio-temporal dynamics of gaze allocation as well as various top-down influences, e.g.
+specific tasks that a user has to perform [1, 7] or social norms that a user ideally adheres to during interactions with humans [8].
+Research on gaze behaviour analysis has revealed that eye movements are closely coordinated with head movements and, as such, that head orientation can be used as a proxy to eye gaze [9, 10, 11, 1].
+Consequently, previous works have mainly focused on exploiting eye-head coordination for predicting eye movements [10, 11, 1] while neglecting coordination between eye and full-body movements.
+Furthermore, most existing works have studied human-object interactions [1, 7] and have neglected the more challenging but for VR/AR applications also more practically useful human-human interactions, in which a person’s eye gaze is influenced by the body movements of the interaction partner.
+To fill this gap, in this work we are the first to report comprehensive analyses of the correlations between gaze direction and full-body movements using four publicly available datasets collected in the real world (MoGaze [12]), in VR (ADT [13]) and AR (GIMO [14] and EgoBody [15]).
+These datasets contain human eye and full-body motion data recorded using a body tracking system and an eye tracker during various daily activities.
+Our analyses reveal that in human-object interactions, e.g.
+pick, place, touch, or hold, eye gaze exhibits strong correlations with full-body motion and eye movements precede body movements.
+In contrast, in human-human interactions, e.g.
+chat, teach, or discuss, a person’s gaze is more closely correlated with the body orientation towards the interaction partner.
+Based on these findings we further present Pose2Gaze – the first method to predict eye movements from human full-body poses during human-object and human-human interactions.
+At the core of our method is a novel learning-based eye-body coordination model that based on a convolutional neural network (CNN) and a spatio-temporal graph convolutional neural network (GCN) to extract features from head directions and full-body poses, respectively.
+A convolutional neural network is also used to predict human eye gaze from these extracted features.
+We compare our method with state-of-the-art gaze prediction methods for three different prediction tasks, i.e.
+generating target eye gaze from past, present, and future body poses, respectively, and show that our method outperforms these baselines by a large margin in terms of mean angular error on the MoGaze ($24.0\%$ improvement), ADT ($10.1\%$), GIMO ($21.3\%$), and EgoBody ($28.6\%$) datasets.
+We also evaluate the effectiveness of our method for a sample downstream task, i.e eye-based activity recognition, and demonstrate that using our gaze prediction method results in significant improvements in recognition accuracy.
+The full source code and trained models are available at zhiminghu.net/hu24_pose2gaze.
+The specific contributions of our work are four-fold:
+  • We provide comprehensive analyses of eye-body coordination in human-object and human-human interactions in the real world, VR, and AR, and show that eye gaze is strongly correlated with full-body motion in human-object interactions and is closely linked with the directions between two bodies in human-human interactions.
+  • We propose Pose2Gaze – a novel eye-body coordination model for gaze prediction that combines a CNN and a spatio-temporal GNN to extract features from head directions and full-body poses.
+  • We report extensive experiments on four public datasets for three different prediction tasks and demonstrate significant performance improvements over several state-of-the-art methods.
+  • We demonstrate our method’s effectiveness in the sample downstream task of eye-based activity recognition.
+II Related Work
+II-A Human Movement Prediction
+Human movement prediction is an important research topic in the areas of virtual reality and augmented reality.
+Some researchers focused on predicting human body movements from speech signals.
 For example, Hasegawa et al.
 used recurrent neural networks (RNNs) to generate natural human full-body poses from perceptual features extracted from the input speech audio[16] while Kucherenko et al.
-employed autoencoder neural networks to learn latent representations for human poses and speech signals and then learned the mappings between the two representations to generate human body motions[27].
+employed autoencoder neural networks to learn latent representations for human poses and speech signals and then learned the mappings between the two representations to predict human body motions[17].
 Other researchers used the text transcripts of speech to generate human poses.
@@
 proposed to use a RNN-based encoder to extract features from speech text and a RNN-based decoder to generate human poses[18] while Bhattacharya et al.
-employed an end-to-end transformer network to generate human movements from the text transcripts of speech[3].
+employed an end-to-end Transformer network to predict human movements from the text transcripts of speech[19].
 In addition, there also exist some works that generated human dance motions from input music and achieved good performances[20, 21].
-While previous works typically focused on generating human body movements, they neglected the generation of human eye gaze, which is significant for human-human[16, 8] and human-computer interactions[9, 44, 23].
-In this work, we focus on generating realistic human eye gaze directly from human full-body poses to fill this gap.
-2.2 Eye-body Coordination
+While previous works typically focused on generating human body movements, they neglected the prediction of human eye gaze, which is significant for human-human[22, 23] and human-computer interactions[24, 4, 25].
+In this work, we focus on predicting human eye movements directly from human full-body poses to fill this gap.
+II-B Eye-body Coordination
 The coordination of human eye and body movements is an important research topic in the areas of cognitive science and human-centred computing.
@@
 head, hand, or torso.
-In contrast, we are the first to study the coordinations of eye and full-body movements simultaneously, which paves the way for generating realistic eye gaze from full-body poses.
-2.3 Eye Gaze Prediction
+In contrast, we are the first to study the coordinations of eye and full-body movements simultaneously, which paves the way for predicting eye gaze from full-body poses.
+II-C Eye Gaze Prediction
 Eye gaze prediction is a popular research topic in the areas of computer vision and human-centred computing.
@@
 Hu et al.
-proposed to use users’ head rotation velocities to predict users’ eye gaze positions in immersive virtual environments and achieved good performances in both free-viewing[21, 20] and task-oriented situations[18, 17].
-However, existing works have only explored the effectiveness of head movements on the task of eye gaze prediction.
-In stark contrast, in this work we focus on the coordination of eye and full-body movements and investigate the effectiveness of full-body movements on generating realistic eye gaze.
-3 Analysis of Eye-body Coordination
-3.1 Gaze and Motion Data
-To study the coordination between human eye gaze and full-body movements, we conducted a comprehensive analysis based on the only public motion capture datasets that also offer eye gaze information: MoGaze [26] and GIMO [51].
-The MoGaze dataset contains eye gaze and full-body poses of $21$ human joints from six people performing everyday pick and place activities in an indoor environment.
-GIMO dataset offers eye gaze and full-body poses of $23$ human joints from $11$ people performing daily activities in various indoor scenes.
-The activities include resting (sitting or laying on objects), interacting with objects (touching, holding, stepping on, reaching to objects), and changing the state of objects (opening, pushing, transferring, throwing, picking up, lifting, connecting, screwing, grabbing, swapping objects).
-3.2 Correlation between Eye Gaze and Body Orientations
-We used the forward directions of human body joints available in the datasets to represent body orientations.
-Specifically, we represented body orientations in MoGaze using the forward directions of head, neck, torso, pelvis, and base and indicated body orientations in GIMO using the directions of head, neck, spine3, spine2, spine1, and pelvis.
-We used the cosine similarity to analyse the correlation between eye gaze and body orientations.
+proposed to use users’ head rotation velocities to predict users’ eye gaze positions in immersive virtual environments and achieved good performances [1, 38].
+However, existing works have only explored the effectiveness of head movements on the task of eye gaze prediction and were limited to human-object interactions.
+In stark contrast, in this work we investigate the effectiveness of full-body movements on predicting human eye gaze in both human-object and human-human interaction activities.
+III Analysis of Eye-body Coordination
+III-A Gaze and Pose Data
+We studied the eye-body coordination from two perspectives, i.e.
+the correlation between gaze direction and body orientations and the correlation between gaze direction and body motions (translational movements of the body).
+We conducted a comprehensive analysis based on four public datasets that contain various human-object and human-human interaction activities collected from the real world (MoGaze [12]), VR environments (ADT [13]), as well as AR scenarios (GIMO [14] and EgoBody [15]).
+MoGaze
+The MoGaze dataset contains 180 minutes of full-body motion capture data collected from seven participants performing everyday pick and place activities in an indoor environment.
+One participant’s eye gaze data is not recorded.
+Therefore, we only used the data from six people to analyse eye-body coordination.
+ADT
+The ADT dataset collects human eye gaze and/or full-body pose data performing various indoor activities in two virtual environments (an apartment and an office environment).
+To analyse eye-body coordination, we used the $34$ sequences that contain both eye gaze and full-body pose data.
+Each sequence lasts for around 2 minutes and the activities include decoration, meal, and work.
+GIMO
+The GIMO dataset records eye gaze and full-body poses from $11$ people performing various daily activities in various indoor scenes.
+The whole dataset contains $215$ sequences and each sequence lasts for around 10 seconds.
+The activities can be classified into three categories, i.e.
+change the state of objects (open, push, transfer, throw, pick up, lift, connect, screw, grab, swap objects), interact with objects (touch, hold, step on, reach to objects), and rest (sit or lay on objects).
+The original dataset does not provide activity labels for each sequence.
+To gain a comprehensive analysis, we manually labelled each sequence as one of the three categories.
+EgoBody
+The EgoBody dataset collects eye gaze and full-body pose data from $36$ people performing diverse human-human interaction activities (between two humans) in 15 indoor scenes.
+The whole dataset contains $125$ sequences and each sequence lasts for around 2 minutes.
+The activities include catch, chat, dance, discuss, learn, perform, and teach.
+III-B Correlation between Eye Gaze and Body Orientations
+TABLE I: The cosine similarities between eye gaze direction and the directions of different body joints in the MoGaze, GIMO, and EgoBody datasets.
+Gaze direction is strongly correlated with body orientations, especially with head direction.
+base pelvis torso neck head
+MoGaze pick 0.64 0.60 0.66 0.84 0.92
+place 0.62 0.58 0.63 0.84 0.92
+GIMO change 0.76 0.86 0.86 0.90 0.93
+interact 0.72 0.82 0.83 0.87 0.93
+rest 0.67 0.82 0.83 0.87 0.92
+EgoBody catch 0.90 0.94 0.94 0.96 0.97
+chat 0.81 0.85 0.87 0.90 0.94
+dance 0.82 0.86 0.87 0.93 0.97
+discuss 0.88 0.88 0.91 0.93 0.94
+learn 0.70 0.75 0.77 0.84 0.89
+perform 0.90 0.92 0.92 0.95 0.97
+teach 0.84 0.84 0.86 0.89 0.93
+Prior works have mainly focused on the correlation between eye gaze and head directions [10, 11].
+To gain a comprehensive understanding of eye-body coordination, we simultaneously analysed the correlation between eye gaze and the orientations of different body joints, e.g.
+neck and torso.
+The ADT dataset does not provide the orientations of the body joints and therefore we only performed analysis on the MoGaze, GIMO, and EgoBody datasets.
+Specifically, we employed the forward directions of the body joints to represent body orientations and used the cosine similarity to analyse the correlation between eye gaze and body orientations.
 Cosine similarity measures the similarity between two non-zero vectors by calculating the cosine of the angle between the vectors and produces a value in the range from $-1$ to $+1$, where $-1$ indicates perfect negative correlation, 00 denotes no correlation, and $+1$ represents perfect positive correlation.
-Table 1 summarises the cosine similarities between eye gaze and body orientations in the MoGaze and GIMO datasets.
-As the table shows, in both datasets eye gaze direction is strongly correlated with the directions of different body joints.
-In particular, eye gaze exhibits very high correlation with head direction, achieving a cosine similarity of $0.92$ in MoGaze and $0.93$ in GIMO.
-We also find that the directions of different body joints have very strong correlations with each other, e.g.
-neck and head directions have a cosine similarity of $0.95$ in MoGaze and $0.92$ in GIMO, indicating that there exists redundant information among these directions.
-Therefore, for simplicity, we suggest to use only head direction to generate human eye gaze rather than using all the body directions.
-Gaze Head Neck Torso Pelvis Base
-Gaze 1 0.92 0.84 0.64 0.59 0.63
-Head - 1 0.95 0.81 0.78 0.81
-Neck - - 1 0.94 0.91 0.93
-Torso - - - 1 0.99 0.99
-Pelvis - - - - 1 0.99
-Base - - - - - 1
-(a) Gaze-body orientations in MoGaze
-Gaze Head Neck Spine3 Spine2 Spine1 Pelvis
-Gaze 1 0.93 0.89 0.84 0.77 0.84 0.72
-Head - 1 0.92 0.84 0.77 0.84 0.71
-Neck - - 1 0.97 0.95 0.97 0.90
-Spine3 - - - 1 0.98 0.99 0.94
-Spine2 - - - - 1 0.97 0.97
-Spine1 - - - - - 1 0.93
-Pelvis - - - - - - 1
-(b) Gaze-body orientations in GIMO
-Table 1: The cosine similarities between eye gaze directions and the directions of different body joints in the MoGaze and GIMO datasets.
-Gaze directions are strongly correlated with body orientations, especially with head directions.
-The directions of different body joints have strong correlations with each other.
-To investigate whether there exist time delays between head direction and eye gaze, we added different time intervals between head and gaze directions and further calculated their cosine similarities.
-We can see from Figure 1 that head direction achieves its highest correlation with eye gaze at $100$ $ms$ in MoGaze and 00 $ms$ in GIMO, indicating that there exists little or no time delay between head and eye movements.
-(a) Gaze-head direction in MoGaze
-(b) Gaze-head direction in GIMO
-Figure 1: The cosine similarities between head and gaze directions at different time intervals in the (a) MoGaze and (b) GIMO datasets.
-Time interval refers to the time difference between head and gaze directions.
-The highest gaze-head correlation occurs at $100$ $ms$ in MoGaze and 00 $ms$ in GIMO, suggesting that there is little or no time delay between head and eye movements.
-3.3 Correlations between Eye Gaze and Body Movements
-To analyse the correlations between human eye gaze and human full-body movements, we calculated the velocities of different body joints and normalised these velocities to 3D unit vectors to represent the directions of body motions.
+Table I summarises the cosine similarities between eye gaze direction and the directions of different body joints in the MoGaze, GIMO, and EgoBody datasets.
+We can see that most of the cosine similarities are larger than $0.6$, demonstrating that eye gaze has very high correlations with body orientations in various daily activities.
+In particular, we find that in all the activities eye gaze consistently exhibits highest correlation with head direction ($>0.9$ in most cases), validating that head direction correlates better with eye gaze than other body orientations.
+To investigate whether there exist time delays between eye gaze and head orientations, we desynchronised the gaze and head signals by adding different time intervals between them (time interval $>0$ means head data is moved forward in time while time interval $<0$ means gaze data is moved forward) and further calculated their cosine similarities.
+We can see from Figure 1 that head direction achieves its highest correlation with eye gaze at the time interval of between $-100$ and $200$ $ms$, indicating that there exists little or no time delay between head and eye movements.
+These results further validate that head direction is a suitable proxy for eye gaze in everyday activities.
+(a) MoGaze
+(b) ADT
+(c) GIMO
+(d) EgoBody
+Figure 1: The cosine similarities between head and gaze directions at different time intervals in the (a) MoGaze (b) ADT (c) GIMO and (d) EgoBody datasets.
+The highest correlations occur at between $-100$ and $200$ $ms$, suggesting that there is little or no time delay between head and eye movements.
+III-C Correlations between Eye Gaze and Body Motions
+To analyse the correlations between human eye gaze and human full-body motion, we calculated the velocities of different body joints using the difference in position between two consecutive frames and then normalised these velocities to 3D unit vectors to represent the directions of body motions.
 We further calculated the cosine similarities between eye gaze and the directions of body motions to measure their correlations.
-Figure 2 shows the cosine similarities between eye gaze and full-body motions in the MoGaze and GIMO datasets, respectively.
-We can see that eye gaze directions have strong correlations with the motions of different body joints in both datasets.
-(a) Gaze-body motions in MoGaze
-(b) Gaze-body motions in GIMO
-Figure 2: The cosine similarities between eye gaze directions and the directions of body motions in the (a) MoGaze and (b) GIMO datasets. Eye gaze is strongly correlated with human full-body motions.
-To investigate any potential time delays between body motions and eye movements, we added different time intervals between body motions and gaze directions and further calculated their correlations.
-Specifically, for simplicity we first grouped all the human joints into three different body parts, i.e.
-Torso, Arm, and Leg.
-In the MoGaze dataset Torso consists of the body joints of base, pelvis, torso, neck, and head while in the GIMO dataset Torso contains the body joints of pelvis, spine1, spine2, spine3, neck, head, and jaw.
-In both datasets, Arm contains the body joints of collar, shoulder, elbow, and wrist while Leg includes hip, knee, ankle, and foot, where the left and right joints are averaged into one single joint, e.g.
-left shoulder and right shoulder are represented using shoulder.
-We then calculated the cosine similarities between eye gaze and the motions of different body parts as well as between eye gaze and the motions of specific joints.
-The correlation between gaze and a body part was computed using the mean of the cosine similarities between eye gaze and the specific joints within the body part.
-We can see from Figure 3 that in both grouped body parts and specific body joints, the highest correlations consistently occur between current eye gaze and body motions in the near future, suggesting that there exists a noticeable time delay between body motions and eye movements.
-(a) Gaze-body motions in MoGaze
-(b) Gaze-torso motions in MoGaze
-(c) Gaze-arm motions in MoGaze
-(d) Gaze-leg motions in MoGaze
-(e) Gaze-body motions in GIMO
-(f) Gaze-torso motions in GIMO
-(g) Gaze-arm motions in GIMO
-(h) Gaze-leg motions in GIMO
-Figure 3: The cosine similarities between eye gaze and the motions of different body parts as well as between eye gaze and the motions of specific body joints at different time intervals in the (a-d) MoGaze and (e-h) GIMO datasets.
-Time interval refers to the time difference between body movements and eye gaze.
-The highest correlations consistently occur between current eye gaze and body motions in the near future (Time Interval $>0$), indicating that there exists a significant time delay between body and eye movements.
-3.4 Summary
-Through comprehensive analysis, we find that eye gaze has very strong correlations with body orientations, especially with head directions, and there exists little or no time delay between head and eye movements.
-We also find that eye gaze is strongly correlated with human full-body motions and there exists a noticeable time delay between body and eye movements.
-These results suggest that head directions and full-body poses contain rich information about eye gaze and thus can be used to generate realistic eye movements.
-4 Pose2Gaze: Eye-body Coordination Model
-We define pose-based eye gaze generation as the task of generating a sequence of eye gaze directions $G_{t+1:t+T}=\{g_{t+1},g_{t+2},...,g_{t+T}\}\in R^{3\times T}$, where $g$ is a 3D unit vector and $T$ is the length of the target eye gaze sequence, from human body orientations and motions.
+Table II shows the cosine similarities between eye gaze and full-body motions in the MoGaze, ADT, GIMO, and EgoBody datasets.
+We find that in the human-object interaction activities, e.g.
+pick, place, decoration, change, and interact, eye gaze exhibits strong correlations with full-body motions ($>$ 0.3 in most cases).
+In the meal and work activities, the eye-body correlations are relatively smaller, probably because the body motions are less frequent in such activities and are thus less correlated with eye gaze.
+In the human-human interaction activities, e.g.
+chat, discuss, and learn, we find that eye gaze has little or no correlation with body motions ($\leq$ 0.05 in most cases).
+This is probably because a person’s eye and body motion are strongly influenced by the interaction partner in human-human interactions, thus degrading the eye-body correlations.
+TABLE II: The cosine similarities between eye gaze and the motions of different body joints in the MoGaze, ADT, GIMO, and EgoBody datasets.
+Eye gaze has strong correlations with body motions in human-object interaction activities while having little or no correlation in human-human interaction activities.
+l_col: left collar, r_col: right collar, l_sho: left shoulder, r_sho: right shoulder, l_elb: left elbow, r_elb: right elbow, l_wri: left wrist, r_wri: right wrist, l_hip: left hip, r_hip: right hip, l_kne: left knee, r_kne: right knee, l_ank: left ankle, r_ank: right ankle, l_toe: left toe, r_toe: right toe.
+base pelvis torso neck head l_col r_col l_sho r_sho l_elb r_elb l_wri r_wri l_hip r_hip l_kne r_kne l_ank r_ank l_toe r_toe Average
+MoGaze pick 0.40 0.40 0.41 0.42 0.46 0.42 0.42 0.38 0.41 0.35 0.40 0.34 0.46 0.40 0.40 0.42 0.42 0.31 0.32 0.37 0.37 0.39
+place 0.48 0.49 0.49 0.50 0.54 0.50 0.50 0.47 0.48 0.44 0.47 0.43 0.58 0.49 0.48 0.50 0.50 0.39 0.39 0.45 0.45 0.48
+ADT decoration 0.28 0.28 0.26 0.26 0.27 0.26 0.26 0.26 0.24 0.27 0.23 0.31 0.25 0.28 0.28 0.28 0.26 0.15 0.12 0.20 0.14 0.25
+meal 0.20 0.20 0.20 0.20 0.20 0.20 0.20 0.20 0.19 0.19 0.20 0.22 0.22 0.20 0.20 0.20 0.20 0.09 0.08 0.13 0.10 0.18
+work 0.18 0.19 0.19 0.20 0.22 0.20 0.20 0.20 0.18 0.19 0.17 0.20 0.18 0.18 0.18 0.19 0.18 0.10 0.09 0.14 0.10 0.17
+GIMO change 0.34 0.34 0.35 0.35 0.34 0.35 0.35 0.34 0.34 0.33 0.33 0.29 0.32 0.34 0.34 0.31 0.31 0.19 0.17 0.15 0.10 0.30
+interact 0.38 0.38 0.37 0.37 0.36 0.37 0.37 0.36 0.36 0.35 0.36 0.32 0.36 0.38 0.38 0.35 0.34 0.21 0.21 0.18 0.15 0.33
+rest 0.36 0.35 0.35 0.34 0.34 0.35 0.35 0.34 0.34 0.32 0.32 0.30 0.32 0.36 0.37 0.33 0.33 0.20 0.18 0.17 0.14 0.31
+EgoBody catch 0.03 0.02 0.02 0.01 0.02 0.02 0.01 0.03 0.00 0.03 -0.02 0.04 0.00 0.03 0.02 0.02 0.02 0.02 0.00 0.02 0.01 0.02
+chat 0.01 0.01 0.01 0.02 0.02 0.01 0.01 0.02 0.02 0.01 0.01 0.01 0.02 0.01 0.01 0.00 0.01 0.01 0.01 0.01 0.01 0.01
+dance 0.05 0.05 0.05 0.04 0.04 0.04 0.05 0.04 0.04 0.03 0.04 0.03 0.03 0.05 0.05 0.05 0.04 0.02 0.01 0.02 0.02 0.04
+discuss 0.02 0.02 0.03 0.03 0.04 0.03 0.03 0.04 0.03 0.02 0.03 0.03 0.03 0.01 0.01 0.00 0.02 0.00 0.00 0.01 0.01 0.02
+learn 0.00 -0.01 -0.01 -0.01 -0.01 -0.01 0.00 -0.01 0.00 -0.01 -0.01 -0.01 0.00 0.00 0.00 -0.01 -0.01 0.00 0.01 0.01 0.01 0.00
+perform 0.04 0.04 0.02 0.02 0.01 0.01 0.03 -0.01 0.01 -0.03 0.01 0.01 0.02 0.04 0.05 0.03 0.01 0.01 0.03 0.02 0.03 0.02
+teach 0.00 0.00 0.01 0.01 0.02 0.01 0.01 0.00 0.02 0.00 0.02 0.01 0.02 0.00 0.01 0.01 0.02 0.02 0.02 0.02 0.02 0.01
+To investigate any potential time delays between body motions and eye movements, we added different time intervals between body motions and gaze directions (time interval $>0$ means body motion data is moved forward in time while time interval $<0$ means gaze data is moved forward) and further calculated their correlations.
+Specifically, we first calculated the cosine similarities between eye gaze and the motions of different body joints and then averaged the cosine similarities across all the body joints.
+We can see from Figure 2 that in human-object interaction activities the highest correlations occur at between $400$ and $1500$ $ms$, suggesting that there exists a noticeable time delay between body motions and eye movements, i.e.
+eye movements precede body motions.
+In the human-human interaction activities, we find that the eye-body correlations are consistently small ($<$ 0.1 in all the cases) at different time intervals, validating that there is little or no correlation between eye gaze and body motion in human-human interactions.
+(a) MoGaze
+(b) ADT
+(c) GIMO
+(d) EgoBody
+Figure 2: The cosine similarities between eye gaze and body motions at different time intervals in the (a) MoGaze (b) ADT (c) GIMO and (d) EgoBody datasets.
+The highest correlations occur at between $400$ and $1500$ $ms$, indicating that eye movements precede body motions.
+III-D Eye-body Coordination in Human-human Interactions
+To analyse the influence of the interaction partner on a person’s eye gaze during human-human interactions, we calculated the directions pointing from a person’s body joints to the body joints of the interaction partner (see Figure 3a).
+To this end we used the difference in position between the person and the interaction partner and normalised these directions to 3D unit vectors.
+We further calculated the cosine similarities between eye gaze and the directions between two bodies to measure their correlation.
+We can see in Table III that eye gaze is highly correlated with the directions between both bodies ($>0.9$ in most cases), suggesting that a person’s eye gaze is significantly influenced by the interaction partner in human-human interactions.
+Figure 3b further shows the cosine similarities between eye gaze and the directions between two bodies at different time intervals.
+We can see that the highest correlations either occur at 00 $ms$ or are very close to the correlation values at 00 $ms$, meaning that there is little or no time delay between body motions and eye movements in human-human interactions.
+TABLE III: The cosine similarities between eye gaze and the directions pointing from a person’s body to the interaction partner in human-human interaction activities.
+Eye gaze is highly correlated with the directions between two bodies.
+l_col: left collar, r_col: right collar, l_sho: left shoulder, r_sho: right shoulder, l_elb: left elbow, r_elb: right elbow, l_wri: left wrist, r_wri: right wrist, l_hip: left hip, r_hip: right hip, l_kne: left knee, r_kne: right knee, l_ank: left ankle, r_ank: right ankle, l_toe: left toe, r_toe: right toe.
+base pelvis torso neck head l_col r_col l_sho r_sho l_elb r_elb l_wri r_wri l_hip r_hip l_kne r_kne l_ank r_ank l_toe r_toe Average
+EgoBody catch 0.92 0.92 0.92 0.92 0.91 0.92 0.92 0.91 0.91 0.90 0.90 0.89 0.89 0.92 0.92 0.91 0.91 0.92 0.91 0.91 0.90 0.91
+chat 0.94 0.94 0.94 0.94 0.94 0.94 0.94 0.92 0.93 0.91 0.91 0.89 0.89 0.93 0.93 0.91 0.92 0.91 0.91 0.89 0.89 0.92
+dance 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.94 0.94 0.91 0.92 0.87 0.90 0.94 0.95 0.92 0.93 0.91 0.93 0.89 0.92 0.93
+discuss 0.93 0.93 0.93 0.93 0.94 0.93 0.93 0.93 0.92 0.92 0.90 0.91 0.88 0.93 0.93 0.92 0.92 0.92 0.92 0.91 0.91 0.92
+learn 0.93 0.92 0.92 0.92 0.92 0.92 0.92 0.92 0.92 0.90 0.91 0.87 0.91 0.92 0.92 0.91 0.92 0.91 0.92 0.89 0.91 0.91
+perform 0.97 0.97 0.97 0.97 0.97 0.97 0.97 0.97 0.96 0.96 0.95 0.95 0.94 0.97 0.97 0.96 0.95 0.96 0.95 0.96 0.94 0.96
+teach 0.93 0.93 0.93 0.93 0.93 0.93 0.93 0.92 0.93 0.91 0.92 0.90 0.91 0.93 0.93 0.92 0.93 0.92 0.92 0.91 0.92 0.92
+Figure 3: (a) Eye gaze and the directions pointing from a person’s body to the body of the interaction partner and (b) the cosine similarities between eye gaze and the directions between two bodies at different time intervals.
+The highest correlations either occur at 00 $ms$ or are very close to the correlation values at 00 $ms$, suggesting that there is little or no time delay between body motions and eye gaze.
+III-E Summary
+Through comprehensive analyses we found that in various daily activities, eye gaze is closely correlated with body orientations, especially with head direction, and there is little or no time delay between head and eye movements.
+In human-object interactions, we find that eye gaze is strongly correlated with human full-body motions and eye movements precede body motions.
+In human-human interactions, we reveal that eye gaze has little or no correlation with body motions while having high correlations with the directions pointing from a person to the interaction partner and that there is little or no time delay between body motions and eye movements.
+These results suggest that head directions and full-body poses contain rich information about eye gaze and thus can be used to predict eye movements.
+IV Pose2Gaze: Eye-body Coordination Model
+IV-A Model Design
+Design of Pose2Gaze
+The analyses in section III revealed that eye gaze is closely correlated with body orientations as well as full-body motions.
+Based on these insights, we propose an eye-body coordination model for gaze prediction that consists of three main components: a body orientation feature extraction module that extracts orientation features, a body motion feature extraction module that extracts motion features, and an eye gaze generation module that generates gaze directions from the extracted body orientation and motion features (see Figure 4 for an overview of our method).
+In the body orientation feature extraction module, we use head direction as input since it has higher correlation with eye gaze than other body orientations (see Table I).
+In the body motion feature extraction module, we use human full-body poses as input in human-object interaction activities since eye gaze is strongly correlated with full-body motions in this situation (see Table II) and employ the full-body poses of both the person and their interaction partner as input in human-human interactions considering that eye gaze has high correlations with the directions pointing from a person to the interaction partner in this setting (see Table III).
+Problem Formulation
+We define pose-based eye gaze prediction as the task of generating a sequence of eye gaze directions $G_{t+1:t+T}=\{g_{t+1},g_{t+2},...,g_{t+T}\}\in R^{3\times T}$, where $g$ is a 3D unit vector and $T$ is the length of the target eye gaze sequence, from human body orientations and motions.
 We use a sequence of head directions $H_{t+1+\Delta t_{h}:t+T+\Delta t_{h}}=\{h_{t+1+\Delta t_{h}},h_{t+2+\Delta t_{h}},...,h_{t+T+\Delta t_{h}}\}\in R^{3\times T}$ to represent body orientations, where $h$ is a 3D unit vector and $\Delta t_{h}$ is the time interval between the input head directions and target eye gaze.
-We employ a sequence of the 3D positions of all human joints $P_{t+1+\Delta t_{p}:t+T+\Delta t_{p}}=\{p_{t+1+\Delta t_{p}},p_{t+2+\Delta t_{p}},...,p_{t+T+\Delta t_{p}}\}\in R^{3\times N\times T}$ to represent body motions, where $N$ is the number of human joints and $\Delta t_{p}$ is the time interval between the input body motions and target eye gaze.
-By setting different time intervals between the input body movements and the target gaze directions, our model can be trained to generate target eye gaze from past, present, or future body poses, respectively (see subsection 5.1).
-Our method consists of three main components: a body orientation feature extraction module that extracts orientation features from head directions, a body motion feature extraction module that extracts motion features from full-body poses, and an eye gaze generation module that generates gaze directions from the extracted body orientation and motion features (see Figure 4 for an overview of our method).
+We employ a sequence of the 3D positions of all human joints $P_{t+1+\Delta t_{p}:t+T+\Delta t_{p}}=\{p_{t+1+\Delta t_{p}},p_{t+2+\Delta t_{p}},...,p_{t+T+\Delta t_{p}}\}\in R^{3\times N\times T}$ to represent body motions, where $N$ is the number of human joints and $\Delta t_{p}$ is the time interval between the input body poses and target eye gaze.
 Figure 4: Architecture of the proposed Pose2Gaze model.
 Pose2Gaze first uses a 1D convolutional neural network to extract body orientation features from head directions, then applies a spatio-temporal graph convolutional neural network to extract the body motion features from human full-body poses, and finally employs a 1D convolutional neural network to generate human eye gaze from the extracted body orientation and motion features.
-4.1 Body Orientation Feature Extraction
-In light of the good performance of 1D convolutional neural networks for processing time series data [20, 18], we employed three 1D CNN layers to extract body orientation features from the sequence of head directions.
+IV-B Body Orientation Feature Extraction
+In light of the good performance of 1D convolutional neural network for processing head movement data [11, 1, 7], we employed three 1D CNN layers to extract body orientation features from the sequence of head directions.
 Specifically, we used two 1D CNN layers, each with $32$ channels and a kernel size of three, to process the head direction sequence $H\in R^{3\times T}$.
@@
 After the two CNN layers, we used a 1D CNN layer with $32$ channels and a kernel size of three, and a Tanh activation function to obtain the body orientation features $f_{ori}\in R^{32\times T}$.
-4.2 Body Motion Feature Extraction
+IV-C Body Motion Feature Extraction
 Given the effectiveness of discrete cosine transform (DCT) for extracting temporal features from human pose data [39, 40], we first employed DCT to encode human pose $P\in R^{3\times N\times T}$ in the temporal domain using DCT matrix $M_{dct}\in R^{T\times T}$:
@@
 where $P_{dct}\in R^{3\times N\times T}$ is the human pose after DCT transform.
-Considering the good performance of graph convolutional neural networks for processing human pose data [33, 35], we propose two GCN blocks, i.e.
+Considering that graph convolutional network outperforms other architectures like Transformer or RNN for processing human pose data [39], we propose two GCN blocks, i.e.
 a start GCN block and a residual GCN block, to extract motion features from the transformed pose data.
@@
 Residual GCN Block
-The residual GCN block contains $m$ GCN components with each component consisting of a temporal GCN that learns the temporal adjacency matrix $A^{T}_{i}\in R^{2T\times 2T}$, a weight matrix $W^{res}_{i}\in R^{16\times 16}$ that extracts the latent features, a spatial GCN that learns the spatial adjacency matrix $A^{S}_{i}\in R^{N\times N}$, a layer normalisation, a Tanh activation function, and a dropout layer with dropout rate $0.3$ to avoid overfitting.
-We set the number of GCN components $m$ to $4$ and added a residual connection for each GCN component to improve the network flow.
-We further cut the output of the residual GCN block in half in the temporal dimension to reduce the feature dimensions and obtained the spatio-temporal human body motion features $f_{mot}\in R^{16\times N\times T}$.
-4.3 Eye Gaze Generation
+The residual GCN block contains $4$ GCN components with each component consisting of a temporal GCN that learns the temporal adjacency matrix $A^{T}_{i}\in R^{2T\times 2T}$, a weight matrix $W^{res}_{i}\in R^{16\times 16}$ that extracts the latent features, a spatial GCN that learns the spatial adjacency matrix $A^{S}_{i}\in R^{N\times N}$, a layer normalisation, a Tanh activation function, and a dropout layer with dropout rate $0.3$ to avoid overfitting.
+We added a residual connection for each GCN component to improve the network flow.
+We further cut the output of the residual GCN block in half in the temporal dimension to reduce the feature dimensions.
+The output of the residual GCN block was converted back to the original representation space using an inverse discrete cosine transform (IDCT)[39].
+A Tanh activation function was applied after the IDCT to obtain the spatio-temporal human body motion features $f_{mot}\in R^{16\times N\times T}$.
+IV-D Eye Gaze Generation
 To generate eye gaze from the extracted body orientation and motion features, we first aggregated the body motion features $f_{mot}$ along the spatial dimension, i.e.
@@
 The generated eye gaze directions were finally normalised to unit vectors: $\hat{G}_{t+1:t+T}=\{\hat{g}_{t+1},\hat{g}_{t+2},...,\hat{g}_{t+T}\}\in R^{3\times T}$.
-4.4 Loss Function
+IV-E Loss Function
 To train our model, we used the mean angular error between the generated eye gaze directions $\hat{G}_{t+1:t+T}$ and the ground truth gaze directions $G_{t+1:t+T}$ as our loss function:
 $\ell=\frac{1}{T}\sum_{j=t+1}^{t+T}\arccos(\hat{g_{j}}\cdot g_{j}).$ (5)
-5 Experiments and Results
-We conducted extensive experiments to evaluate the performance of our gaze generation model.
-Specifically, we compared our model with state-of-the-art gaze prediction methods that estimate eye gaze from head movements on both the MoGaze and GIMO datasets.
+V Experiments and Results
+We conducted extensive experiments to evaluate the performance of our gaze prediction model.
+Specifically, we compared our model with state-of-the-art gaze prediction methods that estimate eye gaze from head movements on the MoGaze, ADT, GIMO, and EgoBody datasets.
 We tested these methods under three different generation settings, i.e.
@@
 We also performed an ablation study to evaluate the effectiveness of each component used in our model.
-5.1 Experimental Settings
+V-A Experimental Settings
 Datasets
-We evaluated our method on both the MoGaze and GIMO datasets since they are the two most established public datasets that offer both eye gaze and human full-body poses.
-MoGaze contains eye gaze and full-body poses recorded at $30$ fps from six people performing everyday pick and place actions.
-To evaluate our model’s generalisation capability for different users, we performed a leave-one-person-out cross-validation: We trained on the data from five people and tested on the remaining one, repeated this procedure six times by testing for a different target person, and calculated the average performance across all six iterations.
-GIMO contains eye gaze and pose data ($30$ fps) from $11$ people performing daily activities in various indoor scenes.
-To evaluate our model on GIMO, we used the default training and test sets provided by the authors [51], i.e.
-we used data from $12$ scenes for training and data from $14$ scenes ($12$ known scenes and two new scenes) for testing.
+We evaluated our method on the MoGaze, ADT, GIMO, and EgoBody datasets (see subsection III-A for the details of these datasets).
+On the MoGaze dataset, we performed a leave-one-person-out cross-validation to evaluate our model’s generalisation capability for different users: We trained on the data from five people and tested on the remaining one, repeated this procedure six times by testing for a different target person, and calculated the average performance across all six iterations.
+On the ADT dataset, we randomly selected $24$ sequences for training and used the remaining $10$ sequences for testing.
+On the GIMO dataset, we used the default training and test sets provided by the authors [14], i.e.
+we used data from $12$ scenes for training ($178$ sequences) and data from $14$ scenes ($12$ known scenes and two new scenes) for testing ($37$ sequences).
+On the EgoBody dataset, we used $82$ sequences for training and $43$ sequences for testing following the default training and test sets provided in the original paper [15].
 Evaluation Metric
@@
 Generation Settings
-We set our model to generate $15$ frames (corresponding to $500$ ms) of eye gaze directions $G_{t+1:t+15}=\{g_{t+1},g_{t+2},...,g_{t+15}\}$.
+We set our model to generate $15$ frames (corresponding to $500$ ms) of eye gaze directions $G_{t+1:t+15}=\{g_{t+1},g_{t+2},...,g_{t+15}\}$ following the settings used in prior work [1, 11].
 We evaluated our method for three different generation tasks: Generating target eye gaze from past, present, and future body poses, respectively:
@@
   • Generating Gaze from Present Poses: We used the body poses and head directions at the present time $P_{t+1:t+15}=\{p_{t+1},p_{t+2},...,p_{t+15}\}$ and $H_{t+1:t+15}=\{h_{t+1},h_{t+2},...,h_{t+15}\}$ as input to generate the corresponding gaze directions.
-This setting corresponds to estimating human eye gaze in real time [21, 25], which is key for a number of applications including gaze-contingent rendering [21, 20], gaze-based interaction [34], and gaze-guided redirected walking [48].
-  • Generating Gaze from Future Poses: subsection 3.3 reveals that eye gaze has highest correlations with body motions in the near future and this implies that using future body poses may improve the performance of gaze generation.
-To evaluate the effectiveness of future body poses, we used the body poses in the future $15$ frames $P_{t+16:t+30}=\{p_{t+16},p_{t+17},...,p_{t+30}\}$ and the head directions at the present time $H_{t+1:t+15}=\{h_{t+1},h_{t+2},...,h_{t+15}\}$ as input to generate the target eye gaze.
-We used real-time head directions because there exists little or no time delay between head and eye movements (subsection 3.2).
-This setting can be seen as an offline processing way of generating realistic eye movements and is particularly important for creating and animating virtual humans [3, 41].
-5.2 Gaze Generation Results
+This setting corresponds to estimating human eye gaze in real time [10, 36], which is key for a number of applications including gaze-contingent rendering [3], gaze-based interaction [44], and gaze-guided redirected walking [6].
+  • Generating Gaze from Future Poses: subsection III-C reveals that in human-object interactions eye gaze has highest correlations with body motions in the near future and this implies that using future body poses may improve the performance of gaze generation.
+Therefore, we used the body poses in the future $15$ frames $P_{t+16:t+30}=\{p_{t+16},p_{t+17},...,p_{t+30}\}$ and the head directions at the present time $H_{t+1:t+15}=\{h_{t+1},h_{t+2},...,h_{t+15}\}$ as input to generate the target eye gaze.
+We used real-time head directions because there exists little or no time delay between head and eye movements (subsection III-B).
+This setting can be seen as an offline processing way of generating eye movements and is particularly important for the applications that rely on offline analysis of gaze data such as virtual content design and optimisation [9, 5].
+V-B Gaze Generation Results
 Generating Gaze from Past Poses
-Table 2 summarises the performances of different methods for generating eye gaze from past poses on both datasets.
-When evaluating on the MoGaze dataset, we calculated the mean angular errors on individual actions (MoGaze-pick, MoGaze-place) as well as on all the actions (MoGaze-all).
-We can see from the table that our method consistently outperforms the state-of-the-art methods on the MoGaze dataset.
-For pick and place actions, our method achieves an average improvement of $18.7\%$ ($14.8^{\circ}$ vs.
-$18.2^{\circ}$) and $27.0\%$ ($11.1^{\circ}$ vs.
-$15.2^{\circ}$) respectively in terms of mean angular error.
-For all actions on MoGaze, our method achieves an average improvement of $22.0\%$ ($13.1^{\circ}$ vs.
-$16.8^{\circ}$) over the state of the art.
-The action labels for some recordings in the GIMO dataset are missing.
-Therefore, when evaluating on GIMO, we directly calculated the mean angular errors on all the actions (GIMO-all) and indicated the results in Table 2.
-We find that our method significantly outperforms the state of the art, achieving an average improvement of $9.7\%$ ($18.6^{\circ}$ vs.
-$20.6^{\circ}$).
-We further performed a paired Wilcoxon signed-rank test to compare the performances of our method with the state-of-the-art methods and validated that the differences between our method and the state of the art are statistically significant ($p<0.01$).
-The above results demonstrate that our method has a strong capability of generating eye gaze from past body poses.
-Ours FixationNet[18] DGaze[20] Head Direction
-MoGaze-pick $\textbf{14.8}^{\circ}$ $\textit{18.2}^{\circ}$ $18.3^{\circ}$ $37.8^{\circ}$
-MoGaze-place $\textbf{11.1}^{\circ}$ $\textit{15.2}^{\circ}$ $15.3^{\circ}$ $34.9^{\circ}$
-MoGaze-all $\textbf{13.1}^{\circ}$ $\textit{16.8}^{\circ}$ $16.9^{\circ}$ $36.4^{\circ}$
-GIMO-all $\textbf{18.6}^{\circ}$ $21.1^{\circ}$ $\textit{20.6}^{\circ}$ $23.1^{\circ}$
-Table 2: Mean angular errors of different methods for generating eye gaze from past poses on the MoGaze and GIMO datasets. Best results are in bold while the second best are in italic.
+Table IV-past summarises the performances of different methods for generating eye gaze from past body poses on the MoGaze, ADT, GIMO, and EgoBody datasets.
+We can see that our method outperforms the state-of-the-art methods in terms of both individual actions as well as all the actions in a dataset, achieving an average improvement of $22.0\%$ ($13.1^{\circ}$ vs.
+$16.8^{\circ}$) on MoGaze, $8.0\%$ ($11.5^{\circ}$ vs.
+$12.5^{\circ}$) on ADT, $11.1\%$ ($18.4^{\circ}$ vs.
+$20.7^{\circ}$) on GIMO, and $28.6\%$ ($13.2^{\circ}$ vs.
+$18.5^{\circ}$) on EgoBody.
+We further performed a paired Wilcoxon signed-rank test to compare the mean angular error of our method and that of the state-of-the-art methods and validated that the differences between our method and the state of the art are statistically significant ($p<0.01$) on all the four datasets.
+In addition, we notice that our method achieves the highest improvement ($28.6\%$) on the EgoBody dataset, demonstrating its effectiveness on the challenging human-human interaction setting.
+We also find that our method’s improvement on the ADT dataset ($8.0\%$) is relatively smaller compared with the improvements on other datasets.
+This is probably because the body motions in the activities of ADT are less frequent and less correlated with eye gaze (see subsection III-C), thus degrading the performance of our method.
+The above results demonstrate that our method has a strong capability of predicting eye gaze from past body poses in both human-object and human-human interaction activities.
+TABLE IV: Mean angular errors of different methods for generating eye gaze from past, present, and future body poses on the MoGaze, ADT, GIMO, and EgoBody datasets.
+Best results are in bold while the second best are underlined.
+MoGaze ADT GIMO EgoBody
+pick place All decoration meal work All change interact rest All catch chat dance discuss learn perform teach All
+past Head Direction $37.8^{\circ}$ $34.9^{\circ}$ $36.4^{\circ}$ $26.5^{\circ}$ $30.6^{\circ}$ $27.1^{\circ}$ $28.0^{\circ}$ $23.5^{\circ}$ $23.7^{\circ}$ $22.9^{\circ}$ $23.4^{\circ}$ $\underline{14.6}^{\circ}$ $18.1^{\circ}$ $25.0^{\circ}$ $18.0^{\circ}$ $17.6^{\circ}$ $\underline{16.8}^{\circ}$ $24.5^{\circ}$ $19.2^{\circ}$
+DGaze[11] $18.3^{\circ}$ $15.3^{\circ}$ $16.9^{\circ}$ $\underline{13.6}^{\circ}$ $\underline{13.2}^{\circ}$ $\underline{11.1}^{\circ}$ $\underline{12.5}^{\circ}$ $23.1^{\circ}$ $\underline{20.4}^{\circ}$ $\underline{18.9}^{\circ}$ $20.9^{\circ}$ $17.1^{\circ}$ $17.9^{\circ}$ $27.1^{\circ}$ $19.6^{\circ}$ $17.3^{\circ}$ $21.0^{\circ}$ $24.6^{\circ}$ $19.5^{\circ}$
+FixationNet[1] $\underline{18.2}^{\circ}$ $\underline{15.2}^{\circ}$ $\underline{16.8}^{\circ}$ $14.8^{\circ}$ $14.3^{\circ}$ $12.0^{\circ}$ $13.5^{\circ}$ $\underline{22.2}^{\circ}$ $\textbf{20.0}^{\circ}$ $19.7^{\circ}$ $\underline{20.7}^{\circ}$ $15.4^{\circ}$ $\underline{17.3}^{\circ}$ $\underline{23.7}^{\circ}$ $\underline{17.6}^{\circ}$ $\underline{16.4}^{\circ}$ $18.9^{\circ}$ $\underline{24.5}^{\circ}$ $\underline{18.5}^{\circ}$
+Ours $\textbf{15.0}^{\circ}$ $\textbf{11.1}^{\circ}$ $\textbf{13.1}^{\circ}$ $\textbf{12.6}^{\circ}$ $\textbf{12.2}^{\circ}$ $\textbf{10.2}^{\circ}$ $\textbf{11.5}^{\circ}$ $\textbf{17.9}^{\circ}$ $21.2^{\circ}$ $\textbf{16.1}^{\circ}$ $\textbf{18.4}^{\circ}$ $\textbf{12.9}^{\circ}$ $\textbf{13.3}^{\circ}$ $\textbf{19.5}^{\circ}$ $\textbf{16.0}^{\circ}$ $\textbf{8.6}^{\circ}$ $\textbf{13.9}^{\circ}$ $\textbf{13.5}^{\circ}$ $\textbf{13.2}^{\circ}$
+present Head Direction $17.6^{\circ}$ $16.2^{\circ}$ $16.9^{\circ}$ $18.5^{\circ}$ $25.3^{\circ}$ $22.9^{\circ}$ $22.3^{\circ}$ $\underline{20.9}^{\circ}$ $19.9^{\circ}$ $18.6^{\circ}$ $19.8^{\circ}$ $\underline{12.4}^{\circ}$ $16.8^{\circ}$ $\underline{19.0}^{\circ}$ $16.6^{\circ}$ $16.6^{\circ}$ $\underline{14.3}^{\circ}$ $23.7^{\circ}$ $17.7^{\circ}$
+DGaze[11] $13.4^{\circ}$ $12.1^{\circ}$ $12.8^{\circ}$ $\underline{10.3}^{\circ}$ $\underline{10.8}^{\circ}$ $\underline{8.8}^{\circ}$ $\underline{9.9}^{\circ}$ $22.6^{\circ}$ $20.5^{\circ}$ $\underline{17.3}^{\circ}$ $20.2^{\circ}$ $14.1^{\circ}$ $16.5^{\circ}$ $22.0^{\circ}$ $16.3^{\circ}$ $\underline{14.8}^{\circ}$ $17.4^{\circ}$ $24.1^{\circ}$ $17.5^{\circ}$
+FixationNet[1] $\underline{13.2}^{\circ}$ $\underline{11.7}^{\circ}$ $\underline{12.5}^{\circ}$ $11.2^{\circ}$ $11.7^{\circ}$ $9.5^{\circ}$ $10.6^{\circ}$ $21.7^{\circ}$ $\underline{19.6}^{\circ}$ $17.5^{\circ}$ $\underline{19.7}^{\circ}$ $13.9^{\circ}$ $\underline{16.3}^{\circ}$ $21.8^{\circ}$ $\underline{16.1}^{\circ}$ $15.1^{\circ}$ $17.2^{\circ}$ $\underline{23.7}^{\circ}$ $\underline{17.3}^{\circ}$
+Ours $\textbf{10.7}^{\circ}$ $\textbf{9.4}^{\circ}$ $\textbf{10.1}^{\circ}$ $\textbf{9.5}^{\circ}$ $\textbf{9.8}^{\circ}$ $\textbf{8.1}^{\circ}$ $\textbf{9.0}^{\circ}$ $\textbf{15.9}^{\circ}$ $\textbf{17.3}^{\circ}$ $\textbf{15.9}^{\circ}$ $\textbf{16.3}^{\circ}$ $\textbf{12.1}^{\circ}$ $\textbf{13.5}^{\circ}$ $\textbf{16.7}^{\circ}$ $\textbf{14.2}^{\circ}$ $\textbf{9.7}^{\circ}$ $\textbf{12.0}^{\circ}$ $\textbf{13.0}^{\circ}$ $\textbf{13.0}^{\circ}$
+future Head Direction $17.6^{\circ}$ $16.2^{\circ}$ $16.9^{\circ}$ $18.5^{\circ}$ $25.3^{\circ}$ $22.9^{\circ}$ $22.3^{\circ}$ $\underline{20.9}^{\circ}$ $19.9^{\circ}$ $18.6^{\circ}$ $19.8^{\circ}$ $\underline{12.4}^{\circ}$ $16.8^{\circ}$ $\underline{19.0}^{\circ}$ $16.6^{\circ}$ $16.6^{\circ}$ $\underline{14.3}^{\circ}$ $23.7^{\circ}$ $17.7^{\circ}$
+DGaze[11] $13.4^{\circ}$ $12.1^{\circ}$ $12.8^{\circ}$ $\underline{10.3}^{\circ}$ $\underline{10.8}^{\circ}$ $\underline{8.8}^{\circ}$ $\underline{9.9}^{\circ}$ $22.6^{\circ}$ $20.5^{\circ}$ $\underline{17.3}^{\circ}$ $20.2^{\circ}$ $14.1^{\circ}$ $16.5^{\circ}$ $22.0^{\circ}$ $16.3^{\circ}$ $\underline{14.8}^{\circ}$ $17.4^{\circ}$ $24.1^{\circ}$ $17.5^{\circ}$
+FixationNet[1] $\underline{13.2}^{\circ}$ $\underline{11.7}^{\circ}$ $\underline{12.5}^{\circ}$ $11.2^{\circ}$ $11.7^{\circ}$ $9.5^{\circ}$ $10.6^{\circ}$ $21.7^{\circ}$ $\underline{19.6}^{\circ}$ $17.5^{\circ}$ $\underline{19.7}^{\circ}$ $13.9^{\circ}$ $\underline{16.3}^{\circ}$ $21.8^{\circ}$ $\underline{16.1}^{\circ}$ $15.1^{\circ}$ $17.2^{\circ}$ $\underline{23.7}^{\circ}$ $\underline{17.3}^{\circ}$
+Ours $\textbf{10.1}^{\circ}$ $\textbf{8.8}^{\circ}$ $\textbf{9.5}^{\circ}$ $\textbf{9.7}^{\circ}$ $\textbf{9.3}^{\circ}$ $\textbf{7.9}^{\circ}$ $\textbf{8.9}^{\circ}$ $\textbf{15.4}^{\circ}$ $\textbf{16.2}^{\circ}$ $\textbf{14.8}^{\circ}$ $\textbf{15.5}^{\circ}$ $\textbf{11.1}^{\circ}$ $\textbf{13.2}^{\circ}$ $\textbf{15.8}^{\circ}$ $\textbf{14.5}^{\circ}$ $\textbf{9.2}^{\circ}$ $\textbf{11.9}^{\circ}$ $\textbf{13.9}^{\circ}$ $\textbf{12.9}^{\circ}$
 Generating Gaze from Present Poses
-Table 3 summarises the mean angular errors of different methods for generating eye gaze from present body poses.
-We can see that on the MoGaze dataset our method significantly outperforms the state of the art on individual actions (MoGaze-pick, MoGaze-place) as well as on all the actions (MoGaze-all).
-Specifically, for pick, place and all actions, our method achieves an average improvement of $20.5\%$ ($10.5^{\circ}$ vs.
-$13.2^{\circ}$), $20.5\%$ ($9.3^{\circ}$ vs.
-$11.7^{\circ}$), and $20.8\%$ ($9.9^{\circ}$ vs.
-$12.5^{\circ}$) respectively in terms of mean angular error.
-On the GIMO dataset, our method outperforms the state of the art with an average improvement of $17.5\%$ ($16.0^{\circ}$ vs.
-$19.4^{\circ}$).
-A paired Wilcoxon signed-rank test was performed to compare the performances of our method with the state-of-the-art methods and the results indicated that the differences between our method and the state of the art are statistically significant ($p<0.01$).
-Figure 5 illustrates the generation results of different methods on the MoGaze and GIMO datasets.
-We can see that our method exhibits better generation performance than other methods (See supplementary video for more results).
-These results demonstrate that our method is able to generate realistic eye gaze from present body poses.
-Ours FixationNet[18] DGaze[20] Head Direction
-MoGaze-pick $\textbf{10.5}^{\circ}$ $\textit{13.2}^{\circ}$ $13.4^{\circ}$ $17.6^{\circ}$
-MoGaze-place $\textbf{9.3}^{\circ}$ $\textit{11.7}^{\circ}$ $12.1^{\circ}$ $16.2^{\circ}$
-MoGaze-all $\textbf{9.9}^{\circ}$ $\textit{12.5}^{\circ}$ $12.8^{\circ}$ $16.9^{\circ}$
-GIMO-all $\textbf{16.0}^{\circ}$ $19.5^{\circ}$ $\textit{19.4}^{\circ}$ $19.4^{\circ}$
-Table 3: Mean angular errors of different methods for generating eye gaze from present poses on the MoGaze and GIMO datasets.
-Figure 5: Results of different methods for generating eye gaze from present poses on MoGaze and GIMO. The green line indicates the ground truth while the blue line represents the generated eye gaze.
+Table IV-present summarises the mean angular errors of different methods for generating eye gaze from present body poses on the MoGaze, ADT, GIMO, and EgoBody datasets.
+We can see that in terms of both individual actions and all the actions in a dataset, our method achieves better performances than the state of the art.
+Specifically, our method achieves an average improvement of $19.2\%$ ($10.1^{\circ}$ vs.
+$12.5^{\circ}$) on MoGaze, $9.1\%$ ($9.0^{\circ}$ vs.
+$9.9^{\circ}$) on ADT, $17.3\%$ ($16.3^{\circ}$ vs.
+$19.7^{\circ}$) on GIMO, and $24.9\%$ ($13.0^{\circ}$ vs.
+$17.3^{\circ}$) on EgoBody.
+A paired Wilcoxon signed-rank test was performed to compare the mean angular error of our method with that of the state-of-the-art methods and the results indicated that the differences between our method and the state of the art are statistically significant ($p<0.01$) on all the four datasets.
+In addition, we find that our method achieves the highest improvement ($24.9\%$) on the EgoBody dataset, demonstrating its superiority for human-human interaction activities.
+Figure 5 illustrates the prediction results of different methods on the MoGaze and GIMO datasets.
+We can see that the eye gaze directions predicted by our method are more close to the ground truth compared with that generated by other methods (see supplementary video for more results).
+These results demonstrate that our method is able to generate eye gaze from present body poses in various daily activities.
+Figure 5: Results of different methods for generating eye gaze from present poses on the MoGaze and GIMO datasets.
+The green line indicates the ground truth while the blue line represents the predicted eye gaze.
 Generating Gaze from Future Poses
-Table 4 summarises the mean angular errors of different methods for generating eye gaze from future body poses.
-We can see that for pick, place and all actions on the MoGaze dataset, our method consistently outperforms the state of the art, achieving an average improvement of $25.8\%$ ($9.8^{\circ}$ vs.
-$13.2^{\circ}$), $26.5\%$ ($8.6^{\circ}$ vs.
-$11.7^{\circ}$), and $26.4\%$ ($9.2^{\circ}$ vs.
-$12.5^{\circ}$) respectively.
-On the GIMO dataset, our method outperforms the state of the art with an average improvement of $21.6\%$ ($15.2^{\circ}$ vs.
-$19.4^{\circ}$) in terms of mean angular error.
-A paired Wilcoxon signed-rank test was conducted and the results validated that the differences between our method and the state of the art are statistically significant ($p<0.01$).
-The above results demonstrate that our method has high performance in generating eye gaze from future body poses.
-Ours FixationNet[18] DGaze[20] Head Direction
-MoGaze-pick $\textbf{9.8}^{\circ}$ $\textit{13.2}^{\circ}$ $13.4^{\circ}$ $17.6^{\circ}$
-MoGaze-place $\textbf{8.6}^{\circ}$ $\textit{11.7}^{\circ}$ $12.1^{\circ}$ $16.2^{\circ}$
-MoGaze-all $\textbf{9.2}^{\circ}$ $\textit{12.5}^{\circ}$ $12.8^{\circ}$ $16.9^{\circ}$
-GIMO-all $\textbf{15.2}^{\circ}$ $19.5^{\circ}$ $\textit{19.4}^{\circ}$ $19.4^{\circ}$
-Table 4: Mean angular errors of different methods for generating eye gaze from future poses on the MoGaze and GIMO datasets.
-5.3 Summary
-The results in above section demonstrate that our method significantly outperforms the state-of-the-art methods for three different generation tasks.
-Furthermore, by comparing our method’s performances under different generation settings, we find that our method achieves its best performance in using future body poses ($9.2^{\circ}$ on MoGaze and $15.2^{\circ}$ on GIMO), followed by using present body poses ($9.9^{\circ}$ on MoGaze and $16.0^{\circ}$ on GIMO) and then using past poses ($13.1^{\circ}$ on MoGaze and $18.6^{\circ}$ on GIMO).
-This result corresponds with our analysis in subsection 3.3 that eye gaze has highest correlations with body motions in the near future and suggests that in offline applications where future pose is available, e.g.
-creating and animating virtual humans [3, 41], using future body poses could generate more realistic eye gaze.
+Table IV-future summarises the mean angular errors of different methods for generating eye gaze from future body poses on the MoGaze, ADT, GIMO, and EgoBody datasets.
+We can see that our method outperforms prior methods in both individual actions and all the actions in a dataset, achieving an average improvement of $24.0\%$ ($9.5^{\circ}$ vs.
+$12.5^{\circ}$) on MoGaze, $10.1\%$ ($8.9^{\circ}$ vs.
+$9.9^{\circ}$) on ADT, $21.3\%$ ($15.5^{\circ}$ vs.
+$19.7^{\circ}$) on GIMO, and $25.4\%$ ($12.9^{\circ}$ vs.
+$17.3^{\circ}$) on EgoBody.
+A paired Wilcoxon signed-rank test was conducted and the results validated that the differences between our method and the state of the art are statistically significant ($p<0.01$) in all the four datasets.
+The above results demonstrate that our method has high performance in generating eye gaze from future body poses in both human-object and human-human interaction activities.
+Summary
+The above results demonstrate that our method significantly outperforms the state-of-the-art methods for three different eye gaze generation tasks.
+Furthermore, we find that in human-object interaction activities, our method achieves significantly higher accuracies using future body poses than using past body poses: $9.5^{\circ}$ vs.
+$13.1^{\circ}$ on MoGaze, $8.9^{\circ}$ vs.
+$11.5^{\circ}$ on ADT, and $15.5^{\circ}$ vs.
+$18.4^{\circ}$ on GIMO.
+These results correspond with our analysis in subsection III-C that eye gaze has highest correlations with body motions in the near future and suggest that in offline applications where future pose is available, e.g.
+virtual content design and optimisation [9, 5], using future body poses could generate more accurate eye gaze directions.
+In human-human interaction activities, we find that our method using past, present, or future body poses achieve similar performances: $13.2^{\circ}$ vs.
+$13.0^{\circ}$ vs.
+$12.9^{\circ}$ on EgoBody.
+This is because the eye-body correlations in human-human interactions are very close at $-500$ $ms$ (past), 00 $ms$ (present) or $500$ $ms$ (future), as illustrated in Figure 3b.
+Model Size and Time Costs
 Our model has smaller size than the state-of-the-art methods, containing only $0.17M$ trainable parameters while DGaze has $0.27M$ parameters and FixationNet contains $0.37M$ parameters.
@@
 We find that our model requires only $35$ ms to train per batch and $3$ ms to test per batch, suggesting that our model is fast enough for practical usage.
-5.4 Ablation Study
+V-C Ablation Study
 We finally performed an ablation study to evaluate the effectiveness of each component used in our model.
-Effectiveness of Our Model Architecture
-To evaluate the effectiveness of our model architecture, we evaluated different ablated versions of it that did not contain the DCT, spatial GCN, temporal GCN, the body poses input, and the head directions input, and re-trained the ablated models.
-Table 5 summarises the performances of different ablated versions of our method for generating eye gaze on the GIMO dataset.
+Specifically, we respectively evaluated different ablated versions of our model that did not contain the DCT, spatial GCN, temporal GCN, the body poses of the person, the body poses of the interaction partner, and the head directions under three generation settings.
+Table V summarises the mean angular errors of different ablated versions of our method on the ADT, GIMO, and EgoBody datasets.
 We can see that our method consistently outperforms the ablated methods under different generation settings and the results are statistically significant (paired Wilcoxon signed-rank test, $p<0.01$), thus underlining the effectiveness of our model architecture.
-One exception is that our method is not as good as the w/o DCT and w/o T-GCN versions in generating eye gaze from past body poses.
-This is probably because past body poses have weaker correlations with eye gaze than present and future body poses (subsection 3.3), and thus the temporal features extracted from past body poses using DCT and temporal GCN are not as effective as that extracted from present and future poses.
-Ours w/o DCT w/o S-GCN w/o T-GCN w/o Poses w/o Head
-Past $18.6^{\circ}$ $\textbf{18.2}^{\circ}$ $19.0^{\circ}$ $18.3^{\circ}$ $22.8^{\circ}$ $23.2^{\circ}$
-Present $\textbf{16.0}^{\circ}$ $16.6^{\circ}$ $17.3^{\circ}$ $16.1^{\circ}$ $20.7^{\circ}$ $20.4^{\circ}$
-Future $\textbf{15.2}^{\circ}$ $16.4^{\circ}$ $17.2^{\circ}$ $16.0^{\circ}$ $20.7^{\circ}$ $19.5^{\circ}$
-Table 5: Mean angular errors of different ablated versions of our method for generating eye gaze under different settings on GIMO.
-Importance of Different Body Parts
-We further evaluated the importance of different body parts for eye gaze generation.
-To this end, we re-trained our model using different body parts as input, i.e.
-only using the Torso part, Arm part, Leg part, as well as different combinations of these.
-Table 6 shows the performances of our method for generating eye gaze from different body parts on the GIMO dataset.
-As can be seen from the table, our method achieves significantly better performances than other versions (paired Wilcoxon signed-rank test, $p<0.01$), validating that each body part used in our model can help improve the performance of eye gaze generation.
-Ours Torso Arm Leg Torso+Arm Torso+Leg Arm+Leg
-Past $\textbf{18.6}^{\circ}$ $20.7^{\circ}$ $19.4^{\circ}$ $19.7^{\circ}$ $19.6^{\circ}$ $19.6^{\circ}$ $18.8^{\circ}$
-Present $\textbf{16.0}^{\circ}$ $17.8^{\circ}$ $17.4^{\circ}$ $18.0^{\circ}$ $16.2^{\circ}$ $17.0^{\circ}$ $16.9^{\circ}$
-Future $\textbf{15.2}^{\circ}$ $18.6^{\circ}$ $16.4^{\circ}$ $18.3^{\circ}$ $16.0^{\circ}$ $16.7^{\circ}$ $16.7^{\circ}$
-Table 6: Mean angular errors of our method for generating eye gaze from different body parts under different settings on GIMO.
-6 Discussion
+In addition, we find that in human-human interactions removing the body poses of the person or the body poses of the interaction partner significantly degrades the performance of our model: EgoBody-past from $13.2^{\circ}$ to $20.6^{\circ}$ or $18.6^{\circ}$, EgoBody-present from $13.0^{\circ}$ to $18.1^{\circ}$ or $17.9^{\circ}$, EgoBody-future from $12.9^{\circ}$ to $17.1^{\circ}$ or $17.9^{\circ}$.
+These results demonstrate the effectiveness of the two body poses on generating eye gaze in human-human interaction activities.
+TABLE V: Mean angular errors of different ablated versions of our method on the ADT, GIMO, and EgoBody datasets.
+Pose refers to the body pose of the person while Pose_I means the body pose of the interaction partner.
+Ours w/o DCT w/o S-GCN w/o T-GCN w/o Pose w/o Pose_I w/o Head
+ADT past $\textbf{11.5}^{\circ}$ $11.7^{\circ}$ $11.8^{\circ}$ $11.9^{\circ}$ $12.2^{\circ}$ $18.2^{\circ}$
+present $\textbf{9.0}^{\circ}$ $9.1^{\circ}$ $9.4^{\circ}$ $9.1^{\circ}$ $9.5^{\circ}$ $17.7^{\circ}$
+future $\textbf{8.9}^{\circ}$ $9.1^{\circ}$ $9.3^{\circ}$ $9.1^{\circ}$ $9.3^{\circ}$ $16.4^{\circ}$
+GIMO past $\textbf{18.4}^{\circ}$ $19.0^{\circ}$ $19.3^{\circ}$ $19.1^{\circ}$ $21.2^{\circ}$ $22.1^{\circ}$
+present $\textbf{16.3}^{\circ}$ $17.3^{\circ}$ $18.1^{\circ}$ $17.3^{\circ}$ $20.8^{\circ}$ $20.9^{\circ}$
+future $\textbf{15.5}^{\circ}$ $16.6^{\circ}$ $18.1^{\circ}$ $16.7^{\circ}$ $20.8^{\circ}$ $18.8^{\circ}$
+EgoBody past $\textbf{13.2}^{\circ}$ $13.5^{\circ}$ $13.4^{\circ}$ $13.4^{\circ}$ $20.6^{\circ}$ $18.6^{\circ}$ $15.1^{\circ}$
+present $\textbf{13.0}^{\circ}$ $13.1^{\circ}$ $14.3^{\circ}$ $13.3^{\circ}$ $18.1^{\circ}$ $17.9^{\circ}$ $14.5^{\circ}$
+future $\textbf{12.9}^{\circ}$ $13.7^{\circ}$ $14.8^{\circ}$ $13.5^{\circ}$ $17.1^{\circ}$ $17.9^{\circ}$ $15.1^{\circ}$
+VI Eye-based Activity Recognition
+Activity recognition is a popular task in the area of human-centred computing and has great relevance for many VR/AR scenarios including adaptive virtual environment design [2], low-latency predictive interfaces [45, 46], or human-aware intelligent systems [47].
+It is well-known that user activities can be recognised directly from human eye gaze [7, 48].
+As such, eye-based activity recognition serves as a particularly relevant sample downstream task to further evaluate our gaze prediction method.
+The higher the activity recognition accuracy, the better the quality of the predicted gaze sequences.
+Datasets
+We only evaluated on the ADT and EgoBody datasets given that they provide activity labels for each recorded sequence and their sequence lengths are sufficient ($>$ 10 s) for training activity recognition methods [7, 49].
+For both datasets, we used the same training and test sets as in section V.
+Activity Recognition Method
+We used the state-of-the-art method EHTask [7] for activity recognition in VR.
+EHTask takes a sequence of eye gaze data as input, extracts eye features using a 1D CNN and a bidirectional gated recurrent unit (GRU), and finally recognises user activities from the eye features using fully-connected layers.
+Procedure
+We first trained EHTask to recognise user activities from the ground truth eye gaze sequences using default parameters.
+On the ADT dataset, we trained EHTask to recognise three activities: decoration, meal, and work.
+On the EgoBody dataset, we trained EHTask for three activities that have the most recordings, i.e.
+chat, learn, and teach.
+Since activity recognition usually requires a long sequence as input [7, 49], we used 300 frames (corresponding to 10 seconds) of ground truth eye gaze data to train EHTask as in the original paper [7].
+At test time, we used the eye gaze sequences generated from different methods as input to EHTask and evaluated its activity recognition performance.
+Since the gaze prediction methods were trained to predict 15 frames of eye gaze while EHTask requires an input of 300 frames, we first segmented the 300 frames of test data into 20 windows (each window contains 15 frames), then predicted eye gaze for each window, and finally concatenated the gaze predictions of all the windows and used them as input to EHTask to recognise user activities.
+Results
+Table VI shows the recognition performances of using the ground truth eye gaze and the eye gaze predicted from different methods on ADT and EgoBody.
+It can be seen in the table that our method achieves higher recognition accuracies than other methods on both ADT ($70.0\%$ vs.
+$67.3\%$) and EgoBody ($60.1\%$ vs.
+$58.2\%$).
+Using a paired Wilcoxon signed-rank test we confirmed that these differences were statistically significant ($p<0.01$).
+We also notice that our method can achieve comparable performance with the ground truth eye gaze in terms of recognition accuracy ($70.0\%$ vs.
+$74.7\%$ on ADT, and $60.1\%$ vs.
+$62.1\%$ on EgoBody).
+These results illustrate the benefit of our method when using the predicted gaze sequences for downstream tasks with high relevance for VR/AR applications, such as eye-based activity recognition.
+TABLE VI: Eye-based activity recognition accuracies of using ground truth eye gaze and the eye gaze predicted from different methods on ADT and EgoBody.
+Best results are in bold while the second best are underlined.
+GT Ours DGaze[11] FixationNet[1] Head Direction Chance
+ADT 74.7% 70.0% 67.3% 66.8% 40.9% 33.3%
+EgoBody 62.1% 60.1% 52.3% 58.2% 50.3% 33.3%
+VII Discussion
 In this work we have made an important step towards understanding the correlation between eye and full-body movements in daily activities and generating eye gaze from full-body poses.
 Eye-body Coordination
-Our analyses on eye-body coordination in daily activities effectively guided the design of our model.
-Specifically, in the analysis of correlation between eye gaze and body orientations, we find that eye gaze has strong correlations with the orientations of different body joints, especially with head directions, and these body orientations are strongly correlated with each other (Table 1).
-Inspired by this, our model extracts body orientation features directly from head directions rather than all the body orientations (subsection 4.1) because head direction itself contains sufficient information for generating eye gaze.
-We also find that there exists little or no time delay between head and eye movements (Figure 1).
-Therefore, in the task of generating gaze from future poses (subsection 5.1), we used head directions at the present time to generate target eye gaze rather than using future head directions.
-In the analysis of correlation between eye gaze and body motions, we find that eye gaze is strongly correlated with full-body motions (Figure 2).
-Therefore, our model extracts features from human full-body poses to generate realistic eye gaze (subsection 4.2).
-We also find that there exists a noticeable time delay between full-body motions and eye gaze (Figure 3).
-Inspired by this, we used future body poses as input to improve the performance of eye gaze generation in situations where future pose information is available (subsection 5.2).
-The good performance of our model validated that our analyses are effective and significant for the design of eye-body coordination models.
-Pose-based Gaze Generation
-Prior methods have typically generated eye gaze from head movements while our method generates eye gaze from full-body poses.
-Our method significantly outperforms prior methods on both the MoGaze and GIMO datasets for three different generation tasks (Table 2, Table 3, Table 4) and the ablation studies validated that each body part contributes to the performance of our method (Table 6).
-These results reveal the significant potential of human full-body poses for generating realistic eye movements and thus open the promising research direction of pose-based gaze generation.
+Our analyses on eye-body coordination in daily activities revealed novel insights for understanding human eye and body movements that we also used to guide the design of our eye-body coordination model.
+Specifically, in the analysis of correlation between gaze and head directions, we found that there exists little or no time delay between head and eye movements (see Figure 1).
+Therefore, in the task of generating gaze from future poses (see subsection V-A), we used head directions at the present time to generate target eye gaze rather than using future head directions.
+In the analysis of correlation between eye gaze and body motions, we found that gaze is strongly correlated with full-body motions in human-object interactions (see subsection III-C) and has high correlations with the directions between two bodies in human-human interactions (see subsection III-D).
+Therefore, our model uses features from the body pose of the person to generate gaze in human-object interactions and employed the features from the body poses of the person and the interaction partner to predict gaze in human-human interactions (see Figure 4).
+We also found that eye movements precede body motions in human-object interactions (see subsection III-C).
+Exploiting this, we used future body poses as input to improve the performance of eye gaze generation in situations where future pose information is available (see subsection V-B).
+The superior performance of our model shows that our analyses are effective and are valuable for the design of future eye-body coordination models.
+Pose-based Gaze Prediction
+Prior methods have typically generated eye gaze from head movements while we predict eye gaze from both head movements and full-body poses.
+This approach proved highly effective and significantly outperforms prior methods in both human-object and human-human interaction activities for three different eye gaze generation tasks (see Table IV).
+An ablation study confirmed that body poses are instrumental for achieving these performance improvements (see Table V).
+Taken together, these results underline the significant potential of human full-body poses for generating eye movements and thus open the promising research direction of pose-based gaze prediction.
 Applications of Our Method
-We evaluated our method for three different generation tasks that are respectively linked to different kinds of applications (subsection 5.1).
-The fact that our method outperforms the state-of-the-art methods by a large margin under all the three different generation settings (Table 2, Table 3, Table 4) demonstrates that our method has great potential for these numerous applications including dynamic event triggering [14], gaze-contingent rendering [21, 20], gaze-based interaction [9], as well as creating and animating virtual humans [3, 41].
-In addition, with the development of motion capture and motion generation techniques [3, 15], it becomes increasingly easy to obtain human full-body poses.
-Our approach could be integrated with such motion capture and motion generation techniques to generate realistic eye and full-body movements simultaneously.
-Limitations
-Despite these advances, we identified several limitations of our work that we plan to address in the future.
-First, we conducted our analysis and experiments using two public indoor datasets that contain eye and full-body movements in various daily activities.
-However, there exist some other interesting activities that are not covered by these datasets, e.g., narration or conversation actions [3].
-No suitable gaze dataset exists yet, so the coordination of eye and full-body movements for these activities still remains to be explored.
-In addition, all activities in our datasets were performed in indoor environments and it remains to be seen whether eye-body coordination as characterised here also applies to outdoor scenarios.
-Furthermore, the activities we studied only cover the interaction between humans and objects and don’t involve other interaction paradigms, e.g.
-interaction between two humans and interaction between a human and a social robot.
-Eye-body coordination in such interaction paradigms may be different and deserves to be studied in future work.
-Future Work
-Besides overcoming the above limitations, many potential avenues of future work exist.
-First, it is well-known that human eye gaze is influenced by both the bottom-up scene content and the top-down tasks.
-Therefore, it will be interesting to analyse how the bottom-up and top-down factors influence the eye-body coordination.
-In addition, incorporating other modalities such as audio, text, or speech into our model may further boost our model’s gaze generation performance.
+In addition to the sample downstream task of eye-based activity recognition (see section VI), we believe that our method has significant potential for numerous VR/AR applications, including dynamic event triggering [2], gaze-contingent rendering [10, 11], gaze-based interaction [24], or virtual content design and optimisation [9, 5].
+In addition, with further advances in the development of motion capture and motion generation techniques [19, 16], it will become increasingly easy to obtain human full-body poses in VR and AR environments.
+In the future, our approach could be integrated with such motion capture and motion generation techniques to generate eye and full-body movements simultaneously.
+Limitations and Future Work
+Despite these advances, our work has several limitations that we plan to address in the future.
+First, we focused on eye-body coordination in different activities and ignored the influence of the visual scene in front of the person.
+It will be interesting to explore whether eye-body coordination during the same activity will be affected by different scene content.
+If so, this will also open a new avenue for research on ways to incorporate visual scene information into the prediction and to see whether this can improve performance further.
+Second, incorporating other modalities such as facial expressions and audio signals in social interactions may also boost our model’s gaze prediction performance.
+In addition, existing datasets that include body pose and eye gaze have all been collected in indoor environments, thus unfortunately limiting our analyses to indoor settings.
+It remains to be seen whether eye-body coordination as characterised here also generalises to outdoor scenarios.
+Furthermore, existing datasets only offer data for interactions between two humans, thus limiting our analyses to dyadic settings.
+It will be interesting to explore the eye-body coordination during other interaction settings that are important for VR/AR applications, e.g.
+interactions between more humans and interactions between a human and a virtual avatar.
 Finally, generating stylistic eye gaze, e.g.
-eye gaze that can convey different emotions [39], is an interesting avenue to extend our work.
-7 Conclusion
-In this work we were the first to explore the challenging task of generating realistic human eye gaze from human full-body poses.
-We first conducted a comprehensive analysis on the coordination of human eye and full-body movements in everyday activities and revealed that eye gaze is strongly correlated with head directions and also full-body motions.
-Based on these insights, we proposed a novel eye-body coordination model to generate eye gaze from head directions and full-body poses that outperformed the state-of-the-art methods by a large margin for three different generation tasks.
-Taken together, our work provides novel insights into eye-body coordination during daily activities and makes an important step towards more holistic, cross-modal generation of virtual agent behaviour.
+eye gaze that can convey different emotions [32], is an interesting avenue for extending our work in the future.
+VIII Conclusion
+In this work we were the first to explore the challenging task of predicting human eye gaze from full-body poses.
+We first conducted a comprehensive analysis on the coordination of human eye and full-body movements in everyday activities and revealed that in human-object interactions eye gaze is strongly correlated with full-body motions while in human-human interactions a person’s gaze direction has strong correlations with the directions pointing from his body to the body of the interaction partner.
+Based on these insights, we proposed a novel eye-body coordination model to predict eye gaze from head directions and full-body poses that outperformed the state-of-the-art methods by a large margin for three different prediction tasks as well as the application of eye-based activity recognition.
+Taken together, our work provides novel insights into eye-body coordination during daily activities and makes an important step towards pose-based gaze prediction.
+Acknowledgement
+This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2075 – 390740016.
+A.
+Bulling was funded by the European Research Council (ERC) under grant agreement 801708.